### v1

# Towards Adversarial Robustness via Debiased High-Confidence Logit Alignment

Abstract
Despite the significant advances that deep neural networks (DNNs) have achieved in various visual tasks, they still exhibit vulnerability to adversarial examples, leading to serious security concerns. Recent adversarial training techniques have utilized inverse adversarial attacks to generate high-confidence examples, aiming to align the distributions of adversarial examples with the high-confidence regions of their corresponding classes. However, in this paper, our investigation reveals that high-confidence outputs under inverse adversarial attacks are correlated with biased feature activation. Specifically, training with inverse adversarial examples causes the model’s attention to shift towards background features, introducing a spurious correlation bias. To address this bias, we propose Debiased High-Confidence Adversarial Training (DHAT), a novel approach that not only aligns the logits of adversarial examples with debiased high-confidence logits obtained from inverse adversarial examples, but also restores the model’s attention to its normal state by enhancing foreground logit orthogonality. Extensive experiments demonstrate that DHAT achieves state-of-the-art performance and exhibits robust generalization capabilities across various vision datasets. Additionally, DHAT can seamlessly integrate with existing advanced adversarial training techniques for improving the performance.
Introduction
Deep neural networks (DNNs) have demonstrated exceptional performance across a range of visual tasks (Zeng et al. 2024; Li et al. 2024a; Gao et al. 2024). However, these networks remain susceptible to adversarial attacks, which present significant security challenges (Szegedy et al. 2014; Fawzi, Fawzi, and Fawzi 2018). Adversarial attacks introduce subtle perturbations that are imperceptible to humans but can significantly disrupt the inference process of DNNs, leading to incorrect predictions. Adversarial training (AT) is widely recognized as one of the most effective methods for defending against these attacks by incorporating adversarial examples into the training process (Goodfellow, Shlens, and Szegedy 2014a; Shafahi et al. 2019).
Recently, numerous advanced AT-based techniques have been proposed to address the vulnerabilities of DNNs (Li and Liu 2023; Subramanian et al. 2024; Huang et al. 2023). Prominent methods in this domain focus on aligning the distributions of adversarial examples with regions corresponding to their true classes, thereby enhancing classification accuracy under adversarial attacks. For example, MART (Wang et al. 2020) and TRADES (Zhang et al. 2019) aim to align the logits of adversarial examples with their corresponding natural examples, promoting consistency in predictions. Expanding on this, UIAT (Dong et al. 2023) and ACR (Cho, Han, and Kim 2023) extend this approach by generating inverse adversarial examples (Salman et al. 2021) that exhibit higher confidence compared to their natural counterparts. This strategy helps alleviate the negative impact of misclassification of natural examples in alignment and directs adversarial examples toward regions with correct high-confidence classifications.
Figure 1: Impact of inverse adversarial examples on model attention. (a) Attention maps generated by Grad-CAM (Selvaraju et al. 2017a) under adversarial and inverse adversarial attacks using the traditional AT model. (b) Comparison of IoU(%) between model attention regions and the true foreground and background regions using UIAT and AT.
Upon investigation, we find that the high-confidence outputs from inverse adversarial examples are correlated with biased feature activation. By visualizing the feature attention maps of the conventional AT model under both adversarial attack and inverse adversarial attack conditions (see Figure 1(a)), we observe that adversarial attacks disrupt the model’s usual attention, leading to incorrect predictions. In contrast, inverse adversarial attacks amplify the model’s attention on features, and stronger inverse attacks can further improve prediction accuracy (Dong et al. 2023). Nevertheless, this amplified attention often causes a shift from the foreground to background features.
To further investigate the impact of inverse adversarial examples on adversarial training, we compared the Intersection over Union (IoU) between the model’s attention maps and the foreground/background regions for the UIAT model (Dong et al. 2023) (trained with inverse adversarial examples) as well as the conventional AT model. As shown in Figure 1(b), increasing the magnitude $\epsilon$ of inverse adversarial attacks enhances the model’s attention to image features. However, this enhanced attention to features introduces a bias, with the UIAT model showing a remarkable shift towards emphasizing background features compared to the conventional AT model. This suggests a compelling perspective: AT model tends to focus on background features when using inverse adversarial examples for training, resulting in biased feature activation. This shift introduces spurious correlation bias, where the model over-relies on background features but neglects foreground features, leading to degraded generalization (Han et al. 2023; Zhang et al. 2024).
Building upon the above findings, we introduce a method named Debiased High-Confidence Adversarial Training (DHAT). DHAT implements two key techniques: Debiased High-Confidence Logit Regularization (DHLR) and Foreground Logit Orthogonal Enhancement (FLOE). DHLR quantifies the bias towards background feature activation and recalibrates the biased high-confidence logits derived from inverse adversarial examples. This regularization aligns the logits of adversarial examples with debiased high-confidence logits to improve adversarial robustness and mitigate spurious correlation bias. FLOE further refines this process by reducing the correlation between high-confidence logits and background features in affine space, which helps restore the model’s attention to its normal state. Extensive experiments demonstrate that DHAT outperforms state-of-the-art techniques across various datasets, providing superior adversarial robustness and generalization. Additionally, our DHAT can be seamlessly integrated with existing advanced adversarial training methods.
Below, we outline our principal contributions:
  • We identify that the adversarial training model shows a biased feature activation under inverse adversarial attacks. Training with inverse adversarial examples causes the model’s attention to shift towards background features, resulting in spurious correlation bias.
  • We introduce a novel method, Debiased High-Confidence Adversarial Training (DHAT). This method aligns the logits of adversarial examples with debiased high-confidence logits corresponding to their classes, effectively mitigating spurious correlation bias.
  • Extensive experiments validate the superior robustness and generalization of our method against adversarial attacks. Additionally, our approach integrates seamlessly with existing advanced adversarial training strategies.
Related Work
Adversarial Attack
Adversarial attacks introduce subtle perturbations $\delta$ into natural examples $x$, causing DNNs to produce incorrect predictions rather than the true class $y$. These attacks are typically formulated as an optimization problem:
$\hat{x}=\underset{||\delta||_{p}\leq\epsilon}{\max}(\mathcal{L}_{attack}(\theta;x+\delta,y)),$ (1)
where $\hat{x}$ represents the adversarial example, $\mathcal{L}_{attack}$ is the adversarial attack loss function, $\theta$ denotes the model parameters, and $\delta$ is constrained to the $p$-norm ball of radius $\epsilon$.
Numerous studies have investigated the vulnerability of DNNs and have proposed various adversarial attack methods. Goodfellow, Shlens, and Szegedy (2014b) introduced the Fast Gradient Sign Method (FGSM), which compute the adversarial perturbation in the direction of the gradient:
$\delta=\eta\cdot\text{sign}(\nabla_{x}\mathcal{L}_{attack}(\theta;x,y)),$ (2)
where $\eta$ denotes the attack step size. Building on this, Madry et al. (2018) proposed Projected Gradient Descent (PGD), which iteratively updates the gradient of the input data to generate more potent adversarial examples, thereby increasing the success rate of attacks:
$\hat{x}^{t+1}=\Pi_{\mathbb{B}(\epsilon)}\{\hat{x}^{t}+\eta\cdot\text{sign}(\nabla_{\hat{x}^{t}}\mathcal{L}_{attack}(\theta;\hat{x}^{t},y))\},$ (3)
where $\Pi_{\mathbb{B}(\epsilon)}$ projects the perturbation onto the $p$-norm ball of radius $\epsilon$. Furthermore, Carlini and Wagner (2017) developed the Carlini-Wagner (C&W) attack, an optimization-based method capable of producing more imperceptible adversarial examples. Croce and Hein (2020b) introduced AutoAttack, an ensemble attack strategy that is user-independent and parameter-free.
Adversarial Defense
Adversarial training (AT) is a widely used strategy for improving model robustness. The core principle of AT is to include adversarial examples in the training process, thereby aligning the model’s outputs towards the true label distribution (Xie et al. 2020; Fowl et al. 2021). Mathematically, conventional AT methods are formulated as a min-max optimization problem:
[…]
where $(x,y)$ denotes a clean image-label pair sampled from the data distribution $\mathcal{D}$, $\delta$ is constrained to the maximum $p$-norm magnitude $\epsilon$, $\mathcal{L}_{AT}$ represents the training strategy function, and $\mathcal{G}_{\theta}(x,\delta)$ denotes the attack generation procedure incorporating the perturbation $\delta$. The outer minimization optimizes the model parameters $\theta$ to minimize the expected loss over adversarial examples, while the inner maximization identifies the perturbations that maximize this loss, thereby enhancing the model’s robustness.
Recent advancements in AT have introduced several techniques to further improve adversarial robustness. Wu, Xia, and Wang (2020) introduced Adversarial Weight Perturbation (AWP), which enhances robustness by flattening the loss landscape with respect to input variations. Kim et al. (2023) proposed Feature Separation and Recalibration (FSR) to capture useful logits from non-robust features. Wei et al. (2023) developed Class-wise Calibrated Fair Adversarial Training (CFA), which adaptively configures training for each class to achieve category-level robust fairness. Li et al. (2024b) proposed SGLR, which uses self-distillation to refine the soft-label distribution logits, thereby calibrating the adversarial training.
Methods similar to our approach include MART (Wang et al. 2020) and UIAT (Dong et al. 2023). These methods use Kullback–Leibler (KL) divergence to align the logits of adversarial examples with high-confidence logits. MART encourages sharing a similar logit distribution between adversarial examples and their corresponding natural examples. UIAT generates inverse adversarial examples (Salman et al. 2021) that exhibit higher confidence compared to natural examples, thereby enhancing the alignment of adversarial examples towards the high-confidence regions of the decision surface. However, our analysis reveals that the high-confidence outputs from inverse adversarial examples are correlated with biased feature activation. Specifically, training with inverse adversarial examples causes model to shift attention from the foreground to background features. Consequently, we focus on mitigating spurious correlation bias in robust model trained with inverse adversarial examples.
Methodology
In this section, we first introduce the basic notation. Then, we describe the Debiased High-Confidence Logit Regularization (DHLR). Next, we present the Foreground Logit Orthogonal Enhancement (FLOE). Finally, we detail the optimization strategy of our proposed method.
Notation
We denote the training dataset as $(X,Y)=\{(x_{i},y_{i})|~{}i\in(1,2,\dots,N)\}$, where $x$ and $y$ denote the natural input examples and their corresponding labels, respectively. $N$ represents the total number of examples in the dataset. Let $\theta$ denote the parameters of a convolutional neural network (CNN) for image classification. The network’s output logits are denoted as $f_{\theta}(\cdot)$, and the final predicted label is given by $\arg\max f_{\theta}(\cdot)$. Inverse adversarial examples $\check{x}$ can be regarded as examples for which the model correctly predicts the label with high-confidence. These examples are generated by applying image perturbations aimed at minimizing the attack loss (Salman et al. 2021):
$\check{x}=\underset{||\delta||_{p}\leq\epsilon}{\min}(\mathcal{L}_{\text{Inv}}(\theta;x-\delta,y)),$ (5)
where $\mathcal{L}_{\text{Inv}}$ denotes the loss function used to generate inverse adversarial attack. The logits of the inverse adversarial example $\check{x}$ and the adversarial example $\hat{x}$ are computed as $\check{z}=f_{\theta}(\check{x})$ and $\hat{z}=f_{\theta}(\hat{x})$, respectively.
Debiased High-Confidence Logit Regularization
To mitigate the spurious correlation bias introduced by adversarial training with inverse adversarial examples, we propose utilizing debiased high-confidence logits for regularization. Specifically, this approach involves separating and quantifying the spurious bias towards background features under inverse adversarial attacks, and then recalibrating the biased high-confidence logits to align with those of adversarial examples, as illustrated in Figure 2.
First, we quantify the model’s biased activation degree towards background features under inverse adversarial attacks. Following the approach proposed by  (Dong et al. 2020), we use Grad-CAM (Selvaraju et al. 2017b) to generate pixel-wise attention maps $M\in[0,1]^{n\times n}$ for the natural example $x$. We separate the background feature $\check{x}_{(S)}$ from the inverse adversarial examples $\check{x}$ based on the $M$:
$[\check{x}_{(S)}]_{(i,j)}=\mathbb{I}_{(M_{i,j}<\omega)}\cdot\check{x}_{(i,j)},$ (6)
where $\mathbb{I}_{(M_{i,j}<\omega)}$ is an indicator function and $\omega$ is a predefined threshold. Next, we compute the logits for the background feature $\check{x}_{(S)}$:
$\check{z}_{(S)}=f_{\theta}(\check{x}_{(S)}).$ (7)
These logits can quantify the biased degree introduced by background feature activation during network inference.
Figure 2: Overview of Debiased High-Confidence Logits Regularization (DHLR). The method first quantifies the degree of biased activation towards background features and recalibrates biased logits, then aligns the logits of adversarial examples with debiased high-confidence logits.
To recalibrate the biased high-confidence logits, we refine the logits of inverse adversarial examples by subtracting the background feature logits:
$\check{z}^{*}=\check{z}-\check{z}_{(S)},$ (8)
where $\check{z}^{*}$ represents the debiased high-confidence logits.
Next, we introduce a regularizer to align the logits of adversarial examples with these debiased high-confidence logits. This regularization encourages the model to focus on foreground features and mitigates the biased activation of background features. Specifically, our Debiased High-Confidence Logit Regularization (DHLR) is defined as:
$\mathcal{L}_{DHLR}(\check{z}^{*},\hat{z})=\mathcal{L}_{KL}(\phi(\check{z}^{*})||\phi(\hat{z})),$ (9)
where $\phi$ denotes the softmax function, $\mathcal{L}_{KL}$ represents the Kullback–Leibler (KL) divergence.
This regularization term $\mathcal{L}_{DHLR}$ promotes consistency between the logits of adversarial examples $\hat{z}$ and the debiased high-confidence logits $\check{z}^{*}$. This alignment mitigates spurious correlation bias, thereby improving adversarial robustness and generalization.
Foreground Logit Orthogonal Enhancement
Post-hoc recalibration of high-confidence logits through DHLR is still insufficient to restore the model’s attention to its normal state when confronted with inverse adversarial examples. Therefore, we further propose a Foreground Logit Orthogonal Enhancement (FLOE) to mitigate this issue.
The primary motivation behind FLOE is to enhance the model’s focus to foreground features under inverse adversarial attacks. FLOE achieves this by minimizing the correlation of the high-confidence logits $\check{z}$ with respect to the background feature logits $\check{z}_{(S)}$, which is implemented by reducing the projection of $\check{z}$ onto $\check{z}_{(S)}$ in the affine space. The optimization objective for FLOE is formulated as:
$\mathcal{L}_{FLOE}(\check{z},\check{z}_{(S)})=|\check{z}-\frac{\check{z}\cdot\check{z}_{(S)}}{|\check{z}_{(S)}|^{2}}\cdot\check{z}_{(S)}|_{p},$ (10)
where $p$ indicates the norm exponent. The term $\mathcal{L}_{FLOE}$ is designed to alleviate biased feature activation under inverse adversarial attacks.
Optimization Strategy
Finally, we integrate the Debiased High-Confidence Logit Regularization (DHLR) and Foreground Logit Orthogonal Enhancement (FLOE) techniques into a unified training framework Debiased High-Confidence Adversarial Training (DHAT). The overall optimization objective for DHAT is formulated as:
$\mathcal{L}_{DHAT}=$ $\mathcal{L}_{AT}(\hat{z},y)+\lambda_{1}\cdot\mathcal{L}_{DHLR}(\check{z}^{*},\hat{z})$ (11)
$+\lambda_{2}\cdot\mathcal{L}_{FLOE}(\check{z},\check{z}_{(S)}),$
where $\mathcal{L}_{AT}(\hat{z},y)$ represents the standard adversarial training loss, and $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters that control the relative contributions of the DHLR and FLOE terms, respectively. Besides, our DHAT can be seamlessly integrated with various advanced adversarial training strategies represented by $\mathcal{L}_{AT}$ to improve the robustness.
Experiments
Experimental Settings
Datasets and models.
We conduct experiments on the CIFAR-10, CIFAR-100 (Krizhevsky, Hinton et al. 2009), and Imagenette (Howard 2019) datasets. The architecture for our experiments include the mainstream models WRN28-10 (Zagoruyko and Komodakis 2016), ResNet-18 (He et al. 2016), ResNet-50 (He et al. 2016), Inception-V3 (Szegedy et al. 2015), and VGG16 (Simonyan and Zisserman 2014). Detailed information about the datasets is provided in the supplementary material.
Training setup.
We employ the Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9 and a weight decay factor of $5\times 10^{-4}$. Each model is trained for 100 epochs, with the learning rate reduced by a factor of 0.1 at the 80th and 90th epochs. For adversarial example generation, we use a fixed 10-iteration attack. The maximum $\ell_{\infty}$-norm of the adversarial perturbation is set to $\epsilon=8/255$, while the inverse adversarial perturbation is set to $\epsilon=4/255$. Both perturbations use a fixed step size of $\alpha=2/255$. Hyperparameters $\lambda_{1}$ and $\lambda_{2}$ are both set to 1.0 across all datasets. Additional training details are available in the supplementary material.
Evaluation setup.
For robustness performance evaluation, we employ PGD (Madry et al. 2018), C&W (Carlini and Wagner 2017), and AutoAttack (AA) (Croce and Hein 2020b) attacks under the $\ell_{\infty}$ norm. AutoAttack comprises APGD-DLR (Croce and Hein 2020b), APGD-CE (Croce and Hein 2020b), FAB (Croce and Hein 2020a), and Square (Andriushchenko et al. 2020). “Clean” denotes natural examples without adversarial perturbations. Furthermore, we introduce the concept of the robust generalization gap (Yin et al. 2023), referred to as the “Robust Gap”. More details are provided in the supplementary material.
[…]
MART (Wang et al. 2020) 82.99 56.25 55.48 55.45 52.26 50.67 9.52
AWP (Wu, Xia, and Wang 2020) 82.67 57.80 57.21 57.07 54.82 51.90 6.90
FSR (Kim et al. 2023) 82.92 56.69 55.94 55.51 53.93 51.74 7.42
CFA (Wei et al. 2023) 84.43 57.87 56.90 56.64 54.60 51.85 10.36
UIAT (Dong et al. 2023) 82.94 58.66 58.12 58.05 54.11 52.17 7.92
SGLR (Li et al. 2024b) 85.76 57.53 56.91 56.66 54.28 52.07 9.38
DHAT (Ours) 83.95 60.49 59.95 59.87 55.27 53.10 3.51
DHAT-AWP (Ours) 83.21 61.61 61.35 61.27 55.71 53.69 4.54
DHAT-CFA (Ours) 84.49 62.67 62.38 62.22 55.95 54.05 6.33
[…]
MART (Wang et al. 2020) 54.69 32.06 31.90 31.88 28.77 27.25 9.96
AWP (Wu, Xia, and Wang 2020) 57.94 34.01 33.75 33.72 30.74 28.90 7.87
FSR (Kim et al. 2023) 57.48 32.93 32.30 32.26 29.16 27.04 7.84
CFA (Wei et al. 2023) 60.92 33.10 32.56 32.41 30.49 28.04 10.47
UIAT (Dong et al. 2023) 57.65 34.27 33.91 33.85 30.97 29.03 11.70
SGLR (Li et al. 2024b) 61.02 33.43 32.98 32.82 30.72 28.50 15.67
DHAT (Ours) 59.14 35.82 35.33 35.02 31.72 30.17 4.24
DHAT-AWP (Ours) 59.30 36.39 36.15 35.98 32.05 30.38 4.85
DHAT-CFA (Ours) 61.54 37.67 37.15 36.99 32.40 30.93 5.93
[…]
MART (Wang et al. 2020) 81.17 59.71 59.23 59.03 57.17 56.18 15.11
AWP (Wu, Xia, and Wang 2020) 83.25 61.32 60.92 60.74 59.39 57.05 25.65
FSR (Kim et al. 2023) 85.86 60.95 60.31 59.70 57.86 56.31 21.95
CFA (Wei et al. 2023) 83.64 60.59 60.18 60.03 58.26 56.70 23.78
UIAT (Dong et al. 2023) 81.35 62.49 62.54 62.51 61.10 57.84 17.87
SGLR (Li et al. 2024b) 84.99 61.48 60.90 60.64 58.39 57.11 22.05
DHAT (Ours) 82.11 64.06 63.72 63.55 62.63 59.09 12.63
DHAT-AWP (Ours) 83.78 64.57 64.34 64.20 63.45 59.47 16.77
DHAT-CFA (Ours) 84.28 64.91 64.92 64.77 63.58 59.95 15.76
Table 1: Comparison of robustness (%) and robust generalization gap (%). The number in bold indicates the best performance.
Experimental Result
In this part, we conducted comprehensive evaluations of our proposed method across various visual datasets and compared it with mainstream adversarial training (AT) methods. Additionally, DHAT can be integrated with advanced AT strategies. The results are shown in Table 1.
Robust performance.
DHAT consistently demonstrates superior robustness across all datasets. For example, on CIFAR-10, DHAT outperforms the second-best method by 1.93% and 1.16% under PGD-10 and C&W attacks, respectively. DHAT also shows the best performance on more complex datasets such as CIFAR-100 and high-resolution Imagenette datasets. Specifically, on Imagenette, DHAT achieves improvements of 1.57% and 1.25% over the second-best method under PGD-10 and AA attacks, respectively. Transfer attack performance comparisons are also reported in the supplementary material.
Combine with other AT methods.
DHAT can be effectively combined with various AT methods to further enhance adversarial performance. Notably, DHAT can be combined with AWP and CFA, denoted as DHAT-AWP and DHAT-CFA, respectively. However, DHAT is not suitable for combining with methods that rely on spurious correlation to improve performance (e.g., FSR, SLGR). In particular, DHAT-CFA achieves superior performance across all datasets. For instance, on Imagenette, DHAT-CFA provides a 2.84% performance improvement under the AA attack compared to the second-best method. Moreover, combining DHAT with advanced AT surpasses the performance of the original baseline methods in terms of both robustness and clean accuracy.
Robust overfitting.
Our method significantly reduces the robust generalization gap across all evaluated datasets, effectively mitigating the robust overfitting phenomenon. For example, on CIFAR-10, DHAT reduces the robust generalization gap by 4.41% and 6.01% compared to UIAT and MART, respectively. This reduction highlights DHAT’s ability to mitigate spurious correlation bias, thereby improving the model’s overall generalization performance. Detailed visualizations of the learning curve are provided in the supplementary material.
Generalization Analysis
In this part, we evaluate the generalizability of our proposed method by comprehensively evaluating its performance across various network architectures and under different magnitudes of adversarial attacks.
Performance with diverse network architecture.
Dataset: CIFAR-10 ResNet-50 VGG-16 Inception-V3
[…]
MART (Wang et al. 2020) 75.05 49.05 4.57 66.72 44.27 3.60 76.05 49.59 5.91
AWP (Wu, Xia, and Wang 2020) 75.59 50.80 3.91 67.14 45.97 3.47 76.74 50.46 4.07
FSR (Kim et al. 2023) 78.38 50.59 3.97 69.59 44.90 3.30 77.28 50.02 4.73
CFA (Wei et al. 2023) 77.88 50.64 4.27 70.45 45.53 4.66 76.70 49.97 4.46
UIAT (Dong et al. 2023) 78.42 51.00 2.99 68.46 45.27 3.57 75.19 51.23 3.70
SGLR (Li et al. 2024b) 80.39 50.55 5.21 70.90 45.04 3.94 80.62 50.73 5.00
DHAT (Ours) 79.94 51.92 1.85 70.01 46.40 1.34 77.85 51.81 1.92
DHAT-AWP (Ours) 76.84 52.05 2.34 67.41 47.21 2.18 77.12 52.08 2.88
DHAT-CFA (Ours) 80.97 52.38 2.09 70.52 47.83 2.21 78.52 52.67 2.44
Table 2: Comparison of robustness (%) and robust generalization gap (%). The number in bold indicates the best performance.
To evaluate the adaptability of our method across different network architectures, we applied DHAT on ResNet-50, VGG-16, and Inception-V3. The performance comparisons are reported in Table 2. The results demonstrate the DHAT’s consistently superior robustness and generalization across a variety of model architectures. For example, DHAT-CFA outperforms the second-best method by 1.83% under AA attacks on ResNet-50. Similarly, on VGG-16, DHAT significantly reduces the robustness generalization gap by 1.96% compared to the second-best method.
Refer to caption Refer to caption
(a) PGD-10 Attack (b) C&W Attack
Figure 3: Comparisons of robustness (%) under varying $\epsilon$ adversarial attacks using ResNet-18 on CIFAR-10.
Different attack magnitude.
Figure 3 illustrates the performance of DHAT under varying magnitudes of PGD-10 and C&W attacks. Our findings reveal that DHAT maintains a significant performance advantage across all tested attack magnitudes. Compared to other methods, DHAT exhibits improved stability and robustness, with minimal performance degradation even under more severe attacks. This consistent performance across varying attack levels underscores the effectiveness of our approach in mitigating the impact of adversarial perturbations. Further details of the attack analysis are provided in the supplementary material.
Ablation Analysis
In this part, we conduct a comprehensive analysis of the impact of individual components and parameters of our model.
Impact of individual components.
[…]
78.03 55.03 48.35 3.84
78.79 56.56 49.47 1.98
78.50 56.22 49.73 2.06
79.28 57.21 50.25 0.81
Table 3: Comparison of robustness (%) and robust generalization gap (%). The number in bold indicates the best performance.
We evaluate the contributions of the individual components of DHAT, specifically Debiased High-Confidence Logit Regularization (DHLR) and Foreground Logit Orthogonal Enhancement (FLOE), as shown in Table 3. The results indicate that both DHLR and FLOE significantly enhance model performance compared to the UIAT method. They enhance clean accuracy and adversarial robustness, while also reducing the gap in robustness generalization. This highlights the effectiveness of DHLR and FLOE in mitigating spurious correlation biases. Notably, the integration of both DHLR and FLOE results in a more substantial overall improvement in model performance.
Refer to caption Refer to caption
(a) $\mathcal{L}_{DHLR}$ (b) $\mathcal{L}_{FLOE}$
Figure 4: Analysis of hyper-parameters $\lambda_{1}$ and $\lambda_{2}$ using ResNet-18 on CIFAR-10. The robust accuracy denotes the accuracy under PGD-10 attack.
Analysis of parameter sensitivity.
Figure 4 illustrates the sensitivity of the optimization parameters $\lambda_{1}$ and $\lambda_{2}$ in DHAT. Our analysis reveals that these parameters influence the trade-off between adversarial robustness and clean accuracy. As $\lambda_{1}$ increases, adversarial robustness improves, while clean accuracy tends to decrease. This occurs because higher values of $\lambda_{1}$ increase the regularization strength, which enhances the alignment of adversarial example logits with debiased high-confidence logits. However, this stronger regularization effect may also lead to a reduced emphasis on clean accuracy, as the model focuses more on handling adversarial perturbations. Conversely, increasing $\lambda_{2}$ generally improves clean accuracy but may adversely affect adversarial robustness. Larger values of $\lambda_{2}$ help the model restore its attention to the normal state, which can result in less effective alignment of adversarial example logits with the debiased high-confidence logits.
Refer to caption Refer to caption
(a) CIFAR-10 (b) Imagenette
Figure 5: Analysis of hyper-parameters $\omega$ using ResNet-18 on CIFAR-10 and Imagenette. The robust accuracy denotes the accuracy under PGD-10 attack.
Analysis of threshold factor for quantifying spurious bias.
Figure 5 illustrates the impact of threshold values $w$ on datasets with varying resolutions. Our analysis reveals that as the threshold value increases, both clean accuracy and robustness initially improve before sharply declining. This initial improvement occurs because higher thresholds help in better separating accurate background features. However, further increasing the threshold eventually leads to the extraction of more foreground features, which causes significant performance degradation. For low-resolution datasets, the optimal threshold tends to be higher compared to high-resolution datasets, as high-resolution images more readily separate background features.
Figure 6: Visualization of feature activation maps using UIAT and our proposed method DHAT, utilizing Grad-CAM under inverse attack ($\epsilon=-8$) and adversarial attack ($\epsilon=8$). Left-to-right: the original image; normal feature activation maps; feature activation maps under adversarial attacks for UIAT and DHAT; and feature activation maps under inverse adversarial attacks for UIAT and DHAT.
Discussion DHAT Effectiveness
In this part, we analyze the mechanisms underlying the effectiveness of our proposed method, Debiased High-Confidence Adversarial Training (DHAT), with a focus on its capacity to mitigate spurious correlation bias.
We begin by visualizing the feature activation maps of models trained with UIAT and DHAT under both adversarial attacks and inverse adversarial attacks. These visualizations, as depicted in Figure 6, highlight notable differences in feature attention maps between the models. Specifically, our proposed method DHAT restores the model’s attention to its normal state. In contrast, UIAT exhibits biased feature activation, particularly under inverse adversarial attack conditions, where the model’s attention shifts toward background features, reducing its attention to foreground features. DHAT’s effectiveness is attributed to its ability to recalibrate biased high-confidence logits during inverse adversarial attacks and reduce the correlation between high-confidence logits and background features in affine space. This ensures that the alignment with inverse adversarial examples is not compromised by spurious correlation bias.
Refer to caption Refer to caption
(a) IoU with True Foreground (b) IoU with True Background
Figure 7: Intersection over Union (IoU) of attention regions with true foreground and background regions. The comparison is conducted among standard adversarial training (AT), UIAT, and DHAT.
Furthermore, we quantitatively evaluate DHAT’s impact on mitigating spurious correlation bias, we compute the Intersection over Union (IoU) between the model’s attention maps and the true foreground and background regions. Figure 7 illustrates the comparison under inverse adversarial attacks using conventional AT, UIAT, and DHAT models. Our results show that DHAT substantially improves the model’s attention to foreground features while effectively alleviating biased feature activation. Specifically, DHAT achieves higher IoU scores with true foreground regions and lower IoU scores with background regions compared to both AT and UIAT. These findings confirm that DHAT effectively mitigates spurious correlation bias.
Conclusion
In this work, we identify a critical issue in adversarial training: AT model exhibits biased feature activation under inverse adversarial attacks. Specifically, training with inverse adversarial examples leads to a shift in the model’s attention toward background features, resulting in spurious correlation bias. To tackle this challenge, we propose a novel method called Debiased High-Confidence Adversarial Training (DHAT). DHAT incorporates two pivotal techniques: Debiased High-Confidence Logit Regularization (DHLR) and Foreground Logit Orthogonal Enhancement (FLOE). DHLR quantifies the bias towards background feature activation and recalibrates biased high-confidence logits to align the logits of adversarial examples with these debiased high-confidence logits. FLOE further restores the model’s attention to its normal state by reducing the correlation between high-confidence logits and background features in affine space. This work highlights DHAT’s effectiveness in enhancing model resilience and presents an adversarial training framework that integrates seamlessly with advanced training strategies.
Supplementary
This supplementary material provides an in-depth examination of the methodologies and results that support the experiments detailed in our study. It aims to offer comprehensive information that validates the findings and ensures reproducibility and transparency.
Experimentation Details
In this part, we provide a comprehensive overview of the experimental setup.
Datasets Information
CIFAR-10 (Krizhevsky, Hinton et al. 2009) comprises 10 classes, with each image having a resolution of 32x32 pixels. This dataset is commonly used for basic image classification tasks and serves as a benchmark for evaluating the performance of various models on a relatively simple dataset.
CIFAR-100 (Krizhevsky, Hinton et al. 2009) contains 100 classes, with each image also having a resolution of 32x32 pixels. The expanded class set in CIFAR-100 increases complexity, enabling a more rigorous evaluation of models in distinguishing among a broader range of objects.
Imagenette (Howard 2019), a subset of ImageNet, includes 10 classes with images at a higher resolution of 320x320 pixels. This dataset is constructed to provide a more realistic and challenging benchmark due to its increased resolution and closer representation of real-world images.
Assessing models on datasets with a larger number of classes and higher resolutions, such as CIFAR-100 and Imagenette, facilitates in-depth analysis of their generalization capabilities. This ensures that models demonstrate effectiveness not only on simpler tasks but also exhibit resilience when confronted with more complex and diverse data.
Training Setup
In the training setup, we employ the Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9 and a weight decay factor of $5\times 10^{-4}$. The initial learning rate is set to 0.1 for CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al. 2009) datasets, while 0.01 for Imagenette (Howard 2019). All models undergo 100 epochs of training, with the learning rate reduced by a factor of 0.1 at the 80th and 90th epochs. The trade-off hyper-parameters $\lambda_{1}$ and $\lambda_{2}$ are set to 1.0 and 1.0. To maintain experimental fairness, all comparative methods undergo identical training strategy, involving the same number of epochs and the same optimizer. For Grad-CAM to quantify the spurious correlation bias, we use a pre-trained ResNet-18 model under simple training, introducing minimal additional training cost. The experiments were conducted on a single NVIDIA Tesla A100.
For adversarial training, we create adversarial examples through a 10-iteration attack where the maximum $\ell_{\infty}$-norm of the adversarial perturbation is limited to $\epsilon=8/255$, using a step size of $\alpha=2/255$. To ensure the reliability of our experiments, we align the inverse adversarial example generation process with UIAT settings (Dong et al. 2023), employing the same loss function and same perturbation constrained by $\epsilon=4/255$.
Notably, our method does not employ label smoothing techniques. This choice is deliberate to avoid potential conflicts with efforts to mitigate spurious correlation bias, as label smoothing can blur class boundaries. Additionally, unlike UIAT, our approach does not utilize momentum terms to stabilize the generation process of inverse adversarial examples, focusing solely on the intrinsic properties of our method.
Attack ($\epsilon=8$) Performance of UIAT / DHAT
Target: VGG-16 Target: WRN28-10 Target: ResNet-18
$\Rightarrow$ ResNet-50 $\Rightarrow$ Inc-V3 $\Rightarrow$ WRN28-10 $\Rightarrow$ ResNet-50 $\Rightarrow$ VGG16 $\Rightarrow$ Inc-v3 $\Rightarrow$ WRN28-10 $\Rightarrow$ VGG16 $\Rightarrow$ Inc-v3
FGSM 63.42/64.83 63.44/64.51 63.73/64.88 77.11/78.73 78.08/79.27 77.84/78.69 73.01/74.80 74.00/75.03 72.43/73.18
PGD-10 53.09/54.46 52.81/53.96 54.17/55.36 61.93/62.87 64.73/65.70 61.40/62.56 58.56/60.30 59.91/60.89 57.82/58.75
PGD-20 52.73/54.25 52.63/53.70 53.80/55.29 61.52/62.68 64.60/65.65 61.19/62.38 58.13/60.04 59.80/60.78 57.48/58.42
PGD-50 52.68/54.25 52.62/53.67 53.77/55.27 61.51/62.65 64.56/65.68 61.17/62.24 58.15/60.03 59.78/60.83 57.45/58.46
C&W 51.90/53.44 51.38/52.61 55.31/56.35 61.05/62.02 63.32/64.40 60.76/61.98 57.73/60.12 57.44/58.74 57.71/59.25
AA 56.31/57.94 56.02/56.98 59.65/60.77 67.18/68.39 71.86/73.00 64.39/65.83 62.36/63.19 65.19/66.07 60.69/61.94
Table 4: Transfer attack accuracy (%) in the single-model transfer scenario. The number in bold indicates the best accuracy.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) FGSM Attack (b) PGD-10 Attack (c) PGD-20 Attack (d) C&W Attack
Figure 8: Comparisons with varying $\epsilon$ values using ResNet-18 on the CIFAR-10. The $x$-axis represents the $\epsilon$ value, while $y$-axis represents the robust accuracy (%).
Evaluation Setup
Our evaluation consists of two main components: robustness performance and robust generalization assessment. To evaluate robustness performance, we use PGD (Madry et al. 2018), C&W (Carlini and Wagner 2017), and AutoAttack (AA) (Croce and Hein 2020b) within the $\ell_{\infty}$-norm. AutoAttack includes several attack methods such as APGD-DLR (Croce and Hein 2020b), APGD-CE (Croce and Hein 2020b), FAB (Croce and Hein 2020a), and Square (Andriushchenko et al. 2020). Adversarial attacks are generated using a step size of $\alpha=2/255$ and a specified maximum $\ell_{\infty}$-norm. Note that the “Clean” indicates natural examples unaffected by adversarial perturbations.
To evaluate robust generalization, we introduce the concept of the robust generalization gap (Yin et al. 2023) denoted as the “Robust Gap” quantifying the difference in robust performance between training and test sets under adversarial attacks. A smaller robust gap indicates improved robust generalization, reflecting reduced vulnerability to robust overfitting in the adversarial training model.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) MART (b) UIAT (c) SGLR (d) DHAT
Figure 9: Two-dimensional t-SNE embeddings of depth features extracted from the penultimate layer of the ResNet-18 model, trained on CIFAR-10 using various methods. The visualization illustrates the distinctiveness and compactness of the learned feature representations for each method.
Performance Under Transfer Attack
In this part, we evaluate our proposed model’s performance under transfer attacks and compare it with the UIAT (Dong et al. 2023). Transfer attacks evaluate a model’s robustness by testing its performance against adversarial examples generated from different source models, which simulate real-world conditions where attackers may use varied strategies. This testing is essential to ensure that a model’s adversarial defenses generalize effectively beyond its training environment and are robust against diverse attack methods. By evaluating the performance under transfer attacks, we ensure that our model is robust not only to attacks generated by its own architecture but also to those from different models, providing a more comprehensive measure of robustness.
Table 4 presents the robust accuracy of various models under transfer attacks with an $\ell_{\infty}$-norm perturbation of $\epsilon=8$. The table compares the performance of the UIAT method with that of our proposed model across different target models and transfer attack types.
Robustness Performance
Our model consistently outperforms UIAT across all transfer adversarial attacks and target models. For example, when the WRN28-10 model is trained with our proposed DHAT and subjected to AA attacks generated from ResNet-50, VGG16, and Inc-V3 source models, the defense success rate improves by 1.21%, 1.14%, and 1.44% compared to UIAT, respectively. This demonstrates the model’s robustness, highlighting its effectiveness not only in specific attack scenarios but also across various model architectures.
Robustness Across Various Attack Types
Our proposed model DHAT demonstrates enhanced robustness, particularly against more challenging attacks such as PGD-50 and C&W. For instance, when using the ResNet-18 as the source model, our method achieves improvements of up to 1.88% and 2.39% under PGD-50 and C&W attacks from WRN28-10, respectively, compared to UIAT. This highlights the model’s superior capability to withstand adversarial perturbations.
Generalization Across Various Source Models
Our model exhibits strong performance across a range of source models, including VGG-16, WRN28-10, and ResNet-18. This indicates that the improved robustness of our model is not limited to specific attacks generated from source architectures but generalizes effectively across different adversarial settings.
Robustness under Various $\epsilon$ Attacks
In this part, we evaluate the robustness of our proposed model under various levels of adversarial perturbations, quantified by different $\epsilon$ values. We compare our model with two baseline methods, MART (Wang et al. 2020) and UIAT (Dong et al. 2023), and the state-of-the-art SGLR method (Li et al. 2024b) across various attack types, including FGSM, PGD-10, PGD-20, and C&W. We aim to demonstrate the superior generalization and robustness of our model under progressively challenging adversarial conditions. Figure 8 shows the accuracy of different methods under various $\epsilon$ values for each attack type.
The comparative analysis of our model with MART, UIAT, and SGLR reveals several key insights:
Robustness Across Various Attack Types
Our model consistently outperforms the baselines and SGLR across all attack types and $\epsilon$ values. This consistent superiority underscores our model’s robust generalization ability, allowing it to maintain high accuracy even under more challenging adversarial attack conditions.
Generalization Against Various Attacks
The results suggest that the defensive mechanisms in our model are highly effective in mitigating the impact of various adversarial attacks (i.e., FGSM, PGD-10, PGD-20, C&W). This robustness is especially evident in scenarios with stronger attacks and higher $\epsilon$ values, where our model shows superior performance compared to existing methods.
Practical Applicability
The enhanced robustness of our model across different $\epsilon$ values and attack types demonstrates its practical applicability in real-world scenarios, where adversarial perturbations can vary in strength and sophistication.
Visualization of Learned Feature Distributions
In this part, we provide a comparative analysis of the feature distributions learned by our proposed method, DHAT, and benchmark it against MART (Wang et al. 2020), UIAT (Dong et al. 2023), and SGLR (Li et al. 2024b). Figure 9 illustrates these feature distributions using t-SNE visualization, with embeddings obtained from the penultimate layer of the ResNet-18 model trained on CIFAR-10.
Our method, DHAT, demonstrates a marked improvement in feature representation. Specifically, the features learned by DHAT are more compact and distinct compared to those produced by the other methods. This enhanced discriminability stems from DHAT’s effectiveness in mitigating spurious correlation bias. By aligning adversarial examples with debiased high-confidence logits, DHAT facilitates the learning of features that are both robust and discriminative, leading to improved performance, particularly in margin-based attacks such as C&W attacks.
When compared between MART and UIAT, MART demonstrates superior feature distinguishability. This advantage stems from MART’s alignment of logit distributions based on natural examples, which promotes the learning of more distinguishable features. In contrast, UIAT aligns the logits of adversarial examples with those of inverse adversarial examples, which are correlated with biased feature activation, resulting in less distinguishable feature representations. Additionally, SGLR’s performance in feature distinguishability is inferior to that of DHAT. This is because SGLR relies on soft-label distributions during training, which tends to rely on spurious correlations, thereby impairing class differentiation.
In summary, DHAT demonstrates superior performance in t-SNE feature distributions, showcasing enhanced feature distinguishability and more effective representations. This improvement is attributed to DHAT’s ability to mitigate spurious correlation bias, which facilitates more robust and accurate feature representations.
Refer to caption Refer to caption
(a) UIAT (b) DHAT
Figure 10: The learning curves for natural accuracy (%) and adversarial accuracy (%) under PGD-10 attack of ResNet-18 on CIFAR-10 dataset.
Robust Overfitting Analysis
In this part, we compare the learning curves of DHAT and UIAT to analyze robust overfitting and generalization. Experiments were conducted using ResNet-18 on the CIFAR-10 dataset, with a total of 200 training epochs for better visualization. Learning rate decay was applied at epochs 100 and 150, reducing the learning rate to 0.1.
Figure 10 shows the learning curves for DHAT and UIAT. Our analysis shows that neither DHAT nor UIAT exhibits a significant or irreversible decline in robustness performance after learning rate adjustments (Yu et al. 2022). This indicates that both methods maintain consistent robustness throughout the training process.
However, DHAT consistently shows a smaller robust generalization gap and standard generalization gap between the training and test sets than UIAT. This reduced performance gap underscores DHAT’s superior generalization ability across different distributions, such as adversarial attack conditions or transitions to unseen test sets. The reduced performance disparity implies that DHAT effectively combats robust overfitting, leading to more reliable and consistent adversarial robustness. In contrast, UIAT shows a more pronounced performance decrease, indicating the degraded generalization capability.
In conclusion, these findings confirm that DHAT not only enhances adversarial robustness but also improves robust generalization, representing a significant advancement in adversarial training techniques.

### v2

@@
 Abstract
-Despite the significant advances that deep neural networks (DNNs) have achieved in various visual tasks, they still exhibit vulnerability to adversarial examples, leading to serious security concerns.
-Recent adversarial training techniques have utilized inverse adversarial attacks to generate high-confidence examples, aiming to align the distributions of adversarial examples with the high-confidence regions of their corresponding classes.
-However, in this paper, our investigation reveals that high-confidence outputs under inverse adversarial attacks are correlated with biased feature activation.
-Specifically, training with inverse adversarial examples causes the model’s attention to shift towards background features, introducing a spurious correlation bias.
-To address this bias, we propose Debiased High-Confidence Adversarial Training (DHAT), a novel approach that not only aligns the logits of adversarial examples with debiased high-confidence logits obtained from inverse adversarial examples, but also restores the model’s attention to its normal state by enhancing foreground logit orthogonality.
-Extensive experiments demonstrate that DHAT achieves state-of-the-art performance and exhibits robust generalization capabilities across various vision datasets.
-Additionally, DHAT can seamlessly integrate with existing advanced adversarial training techniques for improving the performance.
-Introduction
-Deep neural networks (DNNs) have demonstrated exceptional performance across a range of visual tasks (Zeng et al.
-2024; Li et al.
-2024a; Gao et al.
-2024).
-However, these networks remain susceptible to adversarial attacks, which present significant security challenges (Szegedy et al.
-2014; Fawzi, Fawzi, and Fawzi 2018).
+Despite the remarkable progress of deep neural networks (DNNs) in various visual tasks, their vulnerability to adversarial examples raises significant security concerns.
+Recent adversarial training methods leverage inverse adversarial attacks to generate high-confidence examples, aiming to align adversarial distributions with high-confidence class regions.
+However, our investigation reveals that under inverse adversarial attacks, high-confidence outputs are influenced by biased feature activations, causing models to rely on background features that lack a causal relationship with the labels.
+This spurious correlation bias leads to overfitting irrelevant background features during adversarial training, thereby degrading the model’s robust performance and generalization capabilities.
+To address this issue, we propose Debiased High-Confidence Adversarial Training (DHAT), a novel approach that aligns adversarial logits with debiased high-confidence logits and restores proper attention by enhancing foreground logit orthogonality.
+Extensive experiments demonstrate that DHAT achieves state-of-the-art robustness on both CIFAR and ImageNet-1K benchmarks, while significantly improving generalization by mitigating the feature bias inherent in inverse adversarial training approaches.
+Code is available at https://github.com/KejiaZhang-Robust/DHAT.
+1 Introduction
+Deep neural networks (DNNs) have demonstrated exceptional performance across a range of visual tasks [53, 26, 37, 13].
+However, these networks remain susceptible to adversarial attacks, which present significant security challenges [42, 11].
 Adversarial attacks introduce subtle perturbations that are imperceptible to humans but can significantly disrupt the inference process of DNNs, leading to incorrect predictions.
-Adversarial training (AT) is widely recognized as one of the most effective methods for defending against these attacks by incorporating adversarial examples into the training process (Goodfellow, Shlens, and Szegedy 2014a; Shafahi et al.
-2019).
-Recently, numerous advanced AT-based techniques have been proposed to address the vulnerabilities of DNNs (Li and Liu 2023; Subramanian et al.
-2024; Huang et al.
-2023).
-Prominent methods in this domain focus on aligning the distributions of adversarial examples with regions corresponding to their true classes, thereby enhancing classification accuracy under adversarial attacks.
-For example, MART (Wang et al.
-2020) and TRADES (Zhang et al.
-2019) aim to align the logits of adversarial examples with their corresponding natural examples, promoting consistency in predictions.
-Expanding on this, UIAT (Dong et al.
-2023) and ACR (Cho, Han, and Kim 2023) extend this approach by generating inverse adversarial examples (Salman et al.
-2021) that exhibit higher confidence compared to their natural counterparts.
+Adversarial training (AT) is widely recognized as one of the most effective methods for defending against these attacks by incorporating adversarial examples into the training process [14, 36].
+Figure 1: Inverse adversarial attack exploits spurious correlations via attention misallocation to non-causal background features.
+The negative $\epsilon$ denotes inverse adversarial perturbation strength.
+(a) Grad-CAM [34] visualizations show that adversarial attacks disrupt attention and cause errors, while inverse adversarial attacks amplify activation and improve accuracy but expose a critical flaw: attention shifts from foreground to irrelevant background regions.
+(b) Quantitative comparison of IoU(%) between attention regions and the true foreground and background regions using UIAT [9] and AT.
+Increasing inverse attack strength $-\epsilon$ demonstrates UIAT’s disproportionate focus on non-causal background features, exposing reliance on spurious correlations [3, 35, 30].
+Recently, numerous advanced AT-based techniques have been proposed to address the vulnerabilities of DNNs [24, 40, 19, 25].
+Prominent methods in this domain focus on aligning the distributions of adversarial examples to their true classes, thereby enhancing classification accuracy under adversarial attacks.
+For example, MART [45] and TRADES [54] aim to align the logits of adversarial examples with their corresponding natural examples, promoting consistency in predictions.
+Expanding on this, UIAT [9] and ACR [5] extend this approach by generating inverse adversarial examples [32] that exhibit higher confidence compared to their natural counterparts.
 This strategy helps alleviate the negative impact of misclassification of natural examples in alignment and directs adversarial examples toward regions with correct high-confidence classifications.
-Figure 1: Impact of inverse adversarial examples on model attention.
-(a) Attention maps generated by Grad-CAM (Selvaraju et al.
-2017a) under adversarial and inverse adversarial attacks using the traditional AT model.
-(b) Comparison of IoU(%) between model attention regions and the true foreground and background regions using UIAT and AT.
-Upon investigation, we find that the high-confidence outputs from inverse adversarial examples are correlated with biased feature activation.
-By visualizing the feature attention maps of the conventional AT model under both adversarial attack and inverse adversarial attack conditions (see Figure 1(a)), we observe that adversarial attacks disrupt the model’s usual attention, leading to incorrect predictions.
-In contrast, inverse adversarial attacks amplify the model’s attention on features, and stronger inverse attacks can further improve prediction accuracy (Dong et al.
-2023).
-Nevertheless, this amplified attention often causes a shift from the foreground to background features.
-To further investigate the impact of inverse adversarial examples on adversarial training, we compared the Intersection over Union (IoU) between the model’s attention maps and the foreground/background regions for the UIAT model (Dong et al.
-2023) (trained with inverse adversarial examples) as well as the conventional AT model.
-As shown in Figure 1(b), increasing the magnitude $\epsilon$ of inverse adversarial attacks enhances the model’s attention to image features.
-However, this enhanced attention to features introduces a bias, with the UIAT model showing a remarkable shift towards emphasizing background features compared to the conventional AT model.
-This suggests a compelling perspective: AT model tends to focus on background features when using inverse adversarial examples for training, resulting in biased feature activation.
-This shift introduces spurious correlation bias, where the model over-relies on background features but neglects foreground features, leading to degraded generalization (Han et al.
-2023; Zhang et al.
-2024).
+Upon investigation, we uncover a remarkable yet concerning phenomenon: inverse adversarial examples produce high-confidence outputs by triggering biased feature activation.
+To illuminate this effect, we visualize attention maps from a conventional adversarial training (AT) model under both adversarial and inverse adversarial attacks (see Figure 1 (a)).
+Adversarial attacks disrupt normal attention patterns, causing incorrect predictions.
+In contrast, inverse adversarial attacks amplify the model’s attention on features, with stronger attacks even improving prediction accuracy [9, 48].
+However, this apparent enhancement masks a critical flaw: the model’s attention systematically shifts from discriminative foreground features to irrelevant background regions—akin to identifying a sheep by detecting grass rather than the animal itself.
+This reliance on spurious correlations undermines the model’s causal reasoning capabilities and compromises its robustness [20, 6, 1].
+To delve deeper into this issue, we conduct a statistical analysis using the Intersection over Union (IoU) metric to measure the alignment between attention maps and foreground/background regions on a subset of ImageNet [18].
+Leveraging the Class Activation Mapping (CAM)-based technique inspired by [10], we define foreground as regions with high-class activation (i.e., areas critical for predictions) and background as regions with low activation.
+Comparing the UIAT model [9] (trained with inverse adversarial examples) to a standard AT model reveals a concerning trend.
+As evidenced in Figure 1 (b), while increasing the magnitude $\epsilon$ of inverse adversarial attacks enhances overall feature attention, the UIAT model fails to improve IoU for foreground features compared to the AT model.
+Instead, it demonstrates a marked increase in IoU for background features—a clear indicator of attention misallocation.
+This leads to a pivotal insight: the inverse adversarial training disproportionately skews attention toward background features, resulting in biased feature activation.
+This shift engenders a fundamental vulnerability: a spurious correlation bias where the model becomes overly dependent on contextual background cues rather than intrinsic object features.
+Much like a classifier that associates beaches with birds rather than wings and beaks, this misplaced reliance significantly degrades robustness and generalization performance [15, 55], as conclusively demonstrated in Table 5.
 Building upon the above findings, we introduce a method named Debiased High-Confidence Adversarial Training (DHAT).
 DHAT implements two key techniques: Debiased High-Confidence Logit Regularization (DHLR) and Foreground Logit Orthogonal Enhancement (FLOE).
-DHLR quantifies the bias towards background feature activation and recalibrates the biased high-confidence logits derived from inverse adversarial examples.
+DHLR quantifies the bias towards background feature activation and debiase the spurious correlations by recalibrating the biased high-confidence logits derived from inverse adversarial examples.
 This regularization aligns the logits of adversarial examples with debiased high-confidence logits to improve adversarial robustness and mitigate spurious correlation bias.
@@
   • We identify that the adversarial training model shows a biased feature activation under inverse adversarial attacks.
-Training with inverse adversarial examples causes the model’s attention to shift towards background features, resulting in spurious correlation bias.
+Training with inverse adversarial examples causes the model’s attention shifting towards background features, thus lead to spurious correlation bias.
   • We introduce a novel method, Debiased High-Confidence Adversarial Training (DHAT).
@@
   • Extensive experiments validate the superior robustness and generalization of our method against adversarial attacks.
-Additionally, our approach integrates seamlessly with existing advanced adversarial training strategies.
-Related Work
-Adversarial Attack
+Additionally, our approach integrates smoothly with existing advanced adversarial training strategies.
+2 Related Work
+2.1 Adversarial Attack
 Adversarial attacks introduce subtle perturbations $\delta$ into natural examples $x$, causing DNNs to produce incorrect predictions rather than the true class $y$.
@@
 Numerous studies have investigated the vulnerability of DNNs and have proposed various adversarial attack methods.
-Goodfellow, Shlens, and Szegedy (2014b) introduced the Fast Gradient Sign Method (FGSM), which compute the adversarial perturbation in the direction of the gradient:
+Goodfellow et al.
+[14] introduced the Fast Gradient Sign Method (FGSM), which compute the adversarial perturbation in the direction of the gradient:
 $\delta=\eta\cdot\text{sign}(\nabla_{x}\mathcal{L}_{attack}(\theta;x,y)),$ (2)
@@
 Building on this, Madry et al.
-(2018) proposed Projected Gradient Descent (PGD), which iteratively updates the gradient of the input data to generate more potent adversarial examples, thereby increasing the success rate of attacks:
+[29] proposed Projected Gradient Descent (PGD), which iteratively updates the gradient of the input data to generate more potent adversarial examples, thereby increasing the success rate of attacks:
 $\hat{x}^{t+1}=\Pi_{\mathbb{B}(\epsilon)}\{\hat{x}^{t}+\eta\cdot\text{sign}(\nabla_{\hat{x}^{t}}\mathcal{L}_{attack}(\theta;\hat{x}^{t},y))\},$ (3)
 where $\Pi_{\mathbb{B}(\epsilon)}$ projects the perturbation onto the $p$-norm ball of radius $\epsilon$.
-Furthermore, Carlini and Wagner (2017) developed the Carlini-Wagner (C&W) attack, an optimization-based method capable of producing more imperceptible adversarial examples.
-Croce and Hein (2020b) introduced AutoAttack, an ensemble attack strategy that is user-independent and parameter-free.
-Adversarial Defense
+Furthermore, Carlini and Wagner [4] developed the Carlini-Wagner (C&W) attack, an optimization-based method capable of producing more imperceptible adversarial examples.
+Croce and Hein [7] introduced AutoAttack, an ensemble attack strategy that is user-independent and parameter-free.
+2.2 Adversarial Defense
 Adversarial training (AT) is a widely used strategy for improving model robustness.
-The core principle of AT is to include adversarial examples in the training process, thereby aligning the model’s outputs towards the true label distribution (Xie et al.
-2020; Fowl et al.
-2021).
+The core principle of AT is to include adversarial examples in the training process, thereby aligning the model’s outputs towards the true label distribution [50, 12].
 Mathematically, conventional AT methods are formulated as a min-max optimization problem:
@@
 Recent advancements in AT have introduced several techniques to further improve adversarial robustness.
-Wu, Xia, and Wang (2020) introduced Adversarial Weight Perturbation (AWP), which enhances robustness by flattening the loss landscape with respect to input variations.
+Wu et al.
+[49] introduced Adversarial Weight Perturbation (AWP), which enhances robustness by flattening the loss landscape with respect to input variations.
 Kim et al.
-(2023) proposed Feature Separation and Recalibration (FSR) to capture useful logits from non-robust features.
+[21] proposed Feature Separation and Recalibration (FSR) to capture useful logits from non-robust features.
 Wei et al.
-(2023) developed Class-wise Calibrated Fair Adversarial Training (CFA), which adaptively configures training for each class to achieve category-level robust fairness.
+[47] developed Class-wise Calibrated Fair Adversarial Training (CFA), which adaptively configures training for each class to achieve category-level robust fairness.
 Li et al.
-(2024b) proposed SGLR, which uses self-distillation to refine the soft-label distribution logits, thereby calibrating the adversarial training.
-Methods similar to our approach include MART (Wang et al.
-2020) and UIAT (Dong et al.
-2023).
+[27] proposed SGLR, which uses self-distillation to refine the soft-label distribution logits, thereby calibrating the adversarial training.
+Methods similar to our approach include MART [45] and UIAT [9].
 These methods use Kullback–Leibler (KL) divergence to align the logits of adversarial examples with high-confidence logits.
 MART encourages sharing a similar logit distribution between adversarial examples and their corresponding natural examples.
-UIAT generates inverse adversarial examples (Salman et al.
-2021) that exhibit higher confidence compared to natural examples, thereby enhancing the alignment of adversarial examples towards the high-confidence regions of the decision surface.
-However, our analysis reveals that the high-confidence outputs from inverse adversarial examples are correlated with biased feature activation.
-Specifically, training with inverse adversarial examples causes model to shift attention from the foreground to background features.
-Consequently, we focus on mitigating spurious correlation bias in robust model trained with inverse adversarial examples.
-Methodology
-In this section, we first introduce the basic notation.
-Then, we describe the Debiased High-Confidence Logit Regularization (DHLR).
-Next, we present the Foreground Logit Orthogonal Enhancement (FLOE).
-Finally, we detail the optimization strategy of our proposed method.
-Notation
+UIAT generates inverse adversarial examples [32] that exhibit higher confidence compared to natural examples, thereby enhancing the alignment of adversarial examples towards the high-confidence regions of the decision surface.
+3 Methodology
+In this section, we introduce the basic notation and present our approach to mitigate adversarial bias and improve robustness.
+Our method recalibrates high-confidence logits and reduces reliance on spurious background features.
+The optimization strategy is outlined, with pseudocode provided in Section 7 of the supplementary material.
+3.1 Notation
 We denote the training dataset as $(X,Y)=\{(x_{i},y_{i})|~{}i\in(1,2,\dots,N)\}$, where $x$ and $y$ denote the natural input examples and their corresponding labels, respectively.
@@
 Inverse adversarial examples $\check{x}$ can be regarded as examples for which the model correctly predicts the label with high-confidence.
-These examples are generated by applying image perturbations aimed at minimizing the attack loss (Salman et al.
-2021):
+These examples are generated by applying image perturbations aimed at minimizing the attack loss [32]:
 $\check{x}=\underset{||\delta||_{p}\leq\epsilon}{\min}(\mathcal{L}_{\text{Inv}}(\theta;x-\delta,y)),$ (5)
@@
 The logits of the inverse adversarial example $\check{x}$ and the adversarial example $\hat{x}$ are computed as $\check{z}=f_{\theta}(\check{x})$ and $\hat{z}=f_{\theta}(\hat{x})$, respectively.
-Debiased High-Confidence Logit Regularization
+Figure 2: Overview of Debiased High-Confidence Logits Regularization (DHLR).
+The method first quantifies the degree of biased activation towards background features and recalibrates biased logits, then aligns the logits of adversarial examples with debiased high-confidence logits.
+3.2 Debiased High-Confidence Logit Regularization
 To mitigate the spurious correlation bias introduced by adversarial training with inverse adversarial examples, we propose utilizing debiased high-confidence logits for regularization.
@@
 First, we quantify the model’s biased activation degree towards background features under inverse adversarial attacks.
-Following the approach proposed by  (Dong et al.
-2020), we use Grad-CAM (Selvaraju et al.
-2017b) to generate pixel-wise attention maps $M\in[0,1]^{n\times n}$ for the natural example $x$.
-We separate the background feature $\check{x}_{(S)}$ from the inverse adversarial examples $\check{x}$ based on the $M$:
-$[\check{x}_{(S)}]_{(i,j)}=\mathbb{I}_{(M_{i,j}<\omega)}\cdot\check{x}_{(i,j)},$ (6)
-where $\mathbb{I}_{(M_{i,j}<\omega)}$ is an indicator function and $\omega$ is a predefined threshold. Next, we compute the logits for the background feature $\check{x}_{(S)}$:
-$\check{z}_{(S)}=f_{\theta}(\check{x}_{(S)}).$ (7)
+Following the approach proposed by [10], we use using Grad-CAM [33] or alternative methods such as SAM [22] to generate pixel-wise attention maps $M\in[0,1]^{n\times n}$ for the natural example $x$.
+We separate the background feature $\check{x}_{(B)}$ from the inverse adversarial examples $\check{x}$ based on the $M$:
+$[\check{x}_{(B)}]_{(i,j)}=\mathbb{I}_{(M_{i,j}<\omega)}\cdot\check{x}_{(i,j)},$ (6)
+where $\mathbb{I}_{(M_{i,j}<\omega)}$ is an indicator function and $\omega$ is a predefined threshold. Next, we compute the logits for the background feature $\check{x}_{(B)}$:
+$\check{z}_{(B)}=f_{\theta}(\check{x}_{(B)}).$ (7)
 These logits can quantify the biased degree introduced by background feature activation during network inference.
-Figure 2: Overview of Debiased High-Confidence Logits Regularization (DHLR).
-The method first quantifies the degree of biased activation towards background features and recalibrates biased logits, then aligns the logits of adversarial examples with debiased high-confidence logits.
 To recalibrate the biased high-confidence logits, we refine the logits of inverse adversarial examples by subtracting the background feature logits:
-$\check{z}^{*}=\check{z}-\check{z}_{(S)},$ (8)
+$\check{z}^{*}=\check{z}-\check{z}_{(B)},$ (8)
 where $\check{z}^{*}$ represents the debiased high-confidence logits.
@@
 This alignment mitigates spurious correlation bias, thereby improving adversarial robustness and generalization.
-Foreground Logit Orthogonal Enhancement
-Post-hoc recalibration of high-confidence logits through DHLR is still insufficient to restore the model’s attention to its normal state when confronted with inverse adversarial examples.
-Therefore, we further propose a Foreground Logit Orthogonal Enhancement (FLOE) to mitigate this issue.
-The primary motivation behind FLOE is to enhance the model’s focus to foreground features under inverse adversarial attacks.
-FLOE achieves this by minimizing the correlation of the high-confidence logits $\check{z}$ with respect to the background feature logits $\check{z}_{(S)}$, which is implemented by reducing the projection of $\check{z}$ onto $\check{z}_{(S)}$ in the affine space.
-The optimization objective for FLOE is formulated as:
-$\mathcal{L}_{FLOE}(\check{z},\check{z}_{(S)})=|\check{z}-\frac{\check{z}\cdot\check{z}_{(S)}}{|\check{z}_{(S)}|^{2}}\cdot\check{z}_{(S)}|_{p},$ (10)
+3.3 Foreground Logit Orthogonal Enhancement
+While DHLR effectively recalibrates high-confidence logits to mitigate spurious correlations, it fails to fully restore the model’s focus on foreground features when handling inverse adversarial examples $\check{x}$.
+This is because DHLR primarily aligns logits with debiased high-confidence targets, without directly addressing the model’s persistent bias toward background features.
+Under inverse adversarial attacks, the model still exhibits strong activation in irrelevant background regions, undermining its ability to prioritize critical foreground cues.
+To address this issue, we introduce Foreground Logit Orthogonal Enhancement (FLOE), a targeted method designed to minimize the correlation between high-confidence logits $\check{z}$ and background feature logits $\check{z}_{(B)}$.
+The intuition behind FLOE is straightforward: by reducing the projection of $\check{z}$ onto $\check{z}_{(B)}$ in the affine space, we make $\check{z}$ less explainable by $\check{z}_{(B)}$, thereby encouraging the model to rely more on foreground features.
+This projection-based approach ensures that the model’s predictions are driven by meaningful foreground information rather than spurious background correlations.
+The optimization objective is formulated as:
+$\mathcal{L}_{FLOE}(\check{z},\check{z}_{(B)})=-|\check{z}-\frac{\check{z}\cdot\check{z}_{(B)}}{|\check{z}_{(B)}|^{2}}\cdot\check{z}_{(B)}|_{p},$ (10)
 where $p$ indicates the norm exponent. The term $\mathcal{L}_{FLOE}$ is designed to alleviate biased feature activation under inverse adversarial attacks.
-Optimization Strategy
-Finally, we integrate the Debiased High-Confidence Logit Regularization (DHLR) and Foreground Logit Orthogonal Enhancement (FLOE) techniques into a unified training framework Debiased High-Confidence Adversarial Training (DHAT).
-The overall optimization objective for DHAT is formulated as:
+3.4 Optimization Strategy
+Finally, we integrate the DHLR and FLOE techniques into a unified training framework Debiased High-Confidence Adversarial Training (DHAT). The overall optimization objective for DHAT is formulated as:
 $\mathcal{L}_{DHAT}=$ $\mathcal{L}_{AT}(\hat{z},y)+\lambda_{1}\cdot\mathcal{L}_{DHLR}(\check{z}^{*},\hat{z})$ (11)
-$+\lambda_{2}\cdot\mathcal{L}_{FLOE}(\check{z},\check{z}_{(S)}),$
+$+\lambda_{2}\cdot\mathcal{L}_{FLOE}(\check{z},\check{z}_{(B)}),$
 where $\mathcal{L}_{AT}(\hat{z},y)$ represents the standard adversarial training loss, and $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters that control the relative contributions of the DHLR and FLOE terms, respectively.
 Besides, our DHAT can be seamlessly integrated with various advanced adversarial training strategies represented by $\mathcal{L}_{AT}$ to improve the robustness.
-Experiments
-Experimental Settings
-Datasets and models.
-We conduct experiments on the CIFAR-10, CIFAR-100 (Krizhevsky, Hinton et al.
-2009), and Imagenette (Howard 2019) datasets.
-The architecture for our experiments include the mainstream models WRN28-10 (Zagoruyko and Komodakis 2016), ResNet-18 (He et al.
-2016), ResNet-50 (He et al.
-2016), Inception-V3 (Szegedy et al.
-2015), and VGG16 (Simonyan and Zisserman 2014).
-Detailed information about the datasets is provided in the supplementary material.
-Training setup.
-We employ the Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9 and a weight decay factor of $5\times 10^{-4}$.
-Each model is trained for 100 epochs, with the learning rate reduced by a factor of 0.1 at the 80th and 90th epochs.
-For adversarial example generation, we use a fixed 10-iteration attack.
-The maximum $\ell_{\infty}$-norm of the adversarial perturbation is set to $\epsilon=8/255$, while the inverse adversarial perturbation is set to $\epsilon=4/255$.
-Both perturbations use a fixed step size of $\alpha=2/255$.
-Hyperparameters $\lambda_{1}$ and $\lambda_{2}$ are both set to 1.0 across all datasets.
-Additional training details are available in the supplementary material.
-Evaluation setup.
-For robustness performance evaluation, we employ PGD (Madry et al.
-2018), C&W (Carlini and Wagner 2017), and AutoAttack (AA) (Croce and Hein 2020b) attacks under the $\ell_{\infty}$ norm.
-AutoAttack comprises APGD-DLR (Croce and Hein 2020b), APGD-CE (Croce and Hein 2020b), FAB (Croce and Hein 2020a), and Square (Andriushchenko et al.
-2020).
-“Clean” denotes natural examples without adversarial perturbations.
-Furthermore, we introduce the concept of the robust generalization gap (Yin et al.
-2023), referred to as the “Robust Gap”.
-More details are provided in the supplementary material.
+4 Experiments
+Table 1: Comparison of robustness (%) and robust generalization gap (%). The number in bold indicates the best performance.
 […]
 […]
-MART (Wang et al. 2020) 82.99 56.25 55.48 55.45 52.26 50.67 9.52
-AWP (Wu, Xia, and Wang 2020) 82.67 57.80 57.21 57.07 54.82 51.90 6.90
-FSR (Kim et al. 2023) 82.92 56.69 55.94 55.51 53.93 51.74 7.42
-CFA (Wei et al. 2023) 84.43 57.87 56.90 56.64 54.60 51.85 10.36
-UIAT (Dong et al. 2023) 82.94 58.66 58.12 58.05 54.11 52.17 7.92
-SGLR (Li et al. 2024b) 85.76 57.53 56.91 56.66 54.28 52.07 9.38
+MART [45] 62.31 43.56 42.89 42.54 40.62 38.94 13.76
+AWP [49] 64.25 45.13 44.67 44.21 41.96 40.02 12.82
+FSR [21] 64.70 44.31 43.82 43.45 41.23 39.30 14.17
+CFA [47] 65.38 44.87 44.32 43.96 41.58 39.65 15.49
+UIAT [9] 62.64 45.29 44.85 44.57 42.13 40.18 14.68
+SGLR [27] 63.46 44.52 44.05 43.78 41.36 39.47 19.72
+DHAT (Ours) 65.90 46.83 46.42 46.11 43.25 41.70 9.53
+DHAT-AWP (Ours) 65.62 47.38 47.04 46.79 43.62 42.07 10.27
+DHAT-CFA (Ours) 66.26 48.27 47.89 47.53 44.18 42.45 11.64
+[…]
+[…]
+MART [45] 82.99 56.25 55.48 55.45 52.26 50.67 9.52
+AWP [49] 82.67 57.80 57.21 57.07 54.82 51.90 6.90
+FSR [21] 82.92 56.69 55.94 55.51 53.93 51.74 7.42
+CFA [47] 84.43 57.87 56.90 56.64 54.60 51.85 10.36
+UIAT [9] 82.94 58.66 58.12 58.05 54.11 52.17 7.92
+SGLR [27] 85.76 57.53 56.91 56.66 54.28 52.07 9.38
 DHAT (Ours) 83.95 60.49 59.95 59.87 55.27 53.10 3.51
@@
 […]
-MART (Wang et al. 2020) 54.69 32.06 31.90 31.88 28.77 27.25 9.96
-AWP (Wu, Xia, and Wang 2020) 57.94 34.01 33.75 33.72 30.74 28.90 7.87
-FSR (Kim et al. 2023) 57.48 32.93 32.30 32.26 29.16 27.04 7.84
-CFA (Wei et al. 2023) 60.92 33.10 32.56 32.41 30.49 28.04 10.47
-UIAT (Dong et al. 2023) 57.65 34.27 33.91 33.85 30.97 29.03 11.70
-SGLR (Li et al. 2024b) 61.02 33.43 32.98 32.82 30.72 28.50 15.67
+MART [45] 54.69 32.06 31.90 31.88 28.77 27.25 9.96
+AWP [49] 57.94 34.01 33.75 33.72 30.74 28.90 7.87
+FSR [21] 57.48 32.93 32.30 32.26 29.16 27.04 7.84
+CFA [47] 60.92 33.10 32.56 32.41 30.49 28.04 10.47
+UIAT [9] 57.65 34.27 33.91 33.85 30.97 29.03 11.70
+SGLR [27] 61.02 33.43 32.98 32.82 30.72 28.50 15.67
 DHAT (Ours) 59.14 35.82 35.33 35.02 31.72 30.17 4.24
@@
 DHAT-CFA (Ours) 61.54 37.67 37.15 36.99 32.40 30.93 5.93
+4.1 Experimental Settings
+4.1.1 Datasets and models.
+We conduct experiments on the CIFAR-10, CIFAR-100 [23], and ImageNet-1K [31] datasets.
+The architecture for our experiments include the mainstream models WRN28-10 [52], ResNet-18 [16], ResNet-50 [16], Inception-V3 [43], and VGG16 [38].
+Detailed information is provided in the supplementary material.
+4.1.2 Training setup
+For CIFAR datasets, each model is trained for 100 epochs, with the learning rate reduced by a factor of 0.1 at the 80th and 90th epochs.
+For adversarial example generation, we use a fixed 10-iteration attack.
+The maximum $\ell_{\infty}$-norm of the adversarial perturbation is set to $\epsilon=8/255$, while the inverse adversarial perturbation is set to $\epsilon=4/255$.
+Both perturbations use a fixed step size of $\alpha=2/255$.
+For the ImageNet dataset, we adopt the approach in [39] with 2-iteration PGD over 50 epochs, and set $\epsilon=4/255$ and $\alpha=1/255$ for adversarial perturbations.
+The hyperparameters $\lambda_{1}$ and $\lambda_{2}$ are set to 1.0 in all data sets.
+Comprehensive training details are available in the supplementary material.
+4.1.3 Evaluation setup.
+For robustness performance evaluation, we employ PGD [29], C&W [4], and AutoAttack (AA) [7] attacks under the $\ell_{\infty}$ norm.
+AutoAttack comprises APGD-DLR [7], APGD-CE [7], FAB [8], and Square [2].
+“Clean” denotes natural examples without adversarial perturbations.
+Additionally, we introduce the concept of the robust generalization gap, referred to as the “Robust Gap”, which measures the difference in robustness performance between the training and test sets as a direct indicator of the generalization ability [51].
+This metric quantifies the strength of spurious correlations, where larger gaps indicate greater reliance on features that lack causal relationships with target labels.
+Table 2: Comparison of robustness (%) and robust generalization gap (%). The number in bold indicates the best performance.
+Dataset: CIFAR-10 ResNet-50 VGG-16 Inception-V3
 […]
-[…]
-MART (Wang et al. 2020) 81.17 59.71 59.23 59.03 57.17 56.18 15.11
-AWP (Wu, Xia, and Wang 2020) 83.25 61.32 60.92 60.74 59.39 57.05 25.65
-FSR (Kim et al. 2023) 85.86 60.95 60.31 59.70 57.86 56.31 21.95
-CFA (Wei et al. 2023) 83.64 60.59 60.18 60.03 58.26 56.70 23.78
-UIAT (Dong et al. 2023) 81.35 62.49 62.54 62.51 61.10 57.84 17.87
-SGLR (Li et al. 2024b) 84.99 61.48 60.90 60.64 58.39 57.11 22.05
-DHAT (Ours) 82.11 64.06 63.72 63.55 62.63 59.09 12.63
-DHAT-AWP (Ours) 83.78 64.57 64.34 64.20 63.45 59.47 16.77
-DHAT-CFA (Ours) 84.28 64.91 64.92 64.77 63.58 59.95 15.76
-Table 1: Comparison of robustness (%) and robust generalization gap (%). The number in bold indicates the best performance.
-Experimental Result
+MART [45] 75.05 49.05 4.57 66.72 44.27 3.60 76.05 49.59 5.91
+AWP [49] 75.59 50.80 3.91 67.14 45.97 3.47 76.74 50.46 4.07
+FSR [21] 78.38 50.59 3.97 69.59 44.90 3.30 77.28 50.02 4.73
+CFA [47] 77.88 50.64 4.27 70.45 45.53 4.66 76.70 49.97 4.46
+UIAT [9] 78.42 51.00 2.99 68.46 45.27 3.57 75.19 51.23 3.70
+SGLR [27] 80.39 50.55 5.21 70.90 45.04 3.94 80.62 50.73 5.00
+DHAT (Ours) 79.94 51.92 1.85 70.01 46.40 1.34 77.85 51.81 1.92
+DHAT-AWP (Ours) 76.84 52.05 2.34 67.41 47.21 2.18 77.12 52.08 2.88
+DHAT-CFA (Ours) 80.97 52.38 2.09 70.52 47.83 2.21 78.52 52.67 2.44
+4.2 Experimental Result
 In this part, we conducted comprehensive evaluations of our proposed method across various visual datasets and compared it with mainstream adversarial training (AT) methods.
@@
 The results are shown in Table 1.
-Robust performance.
+4.2.1 Robust performance.
 DHAT consistently demonstrates superior robustness across all datasets.
@@
 DHAT also shows the best performance on more complex datasets such as CIFAR-100 and high-resolution Imagenette datasets.
-Specifically, on Imagenette, DHAT achieves improvements of 1.57% and 1.25% over the second-best method under PGD-10 and AA attacks, respectively.
-Transfer attack performance comparisons are also reported in the supplementary material.
-Combine with other AT methods.
+Specifically, on ImageNet-1K, DHAT achieves improvements of 1.54% and 1.52% over the second-best method under PGD-10 and AA attacks, respectively.
+4.2.2 Combine with other AT methods.
 DHAT can be effectively combined with various AT methods to further enhance adversarial performance.
@@
 Moreover, combining DHAT with advanced AT surpasses the performance of the original baseline methods in terms of both robustness and clean accuracy.
-Robust overfitting.
+4.2.3 Robust overfitting.
 Our method significantly reduces the robust generalization gap across all evaluated datasets, effectively mitigating the robust overfitting phenomenon.
@@
 Detailed visualizations of the learning curve are provided in the supplementary material.
-Generalization Analysis
+4.3 Generalization Analysis
 In this part, we evaluate the generalizability of our proposed method by comprehensively evaluating its performance across various network architectures and under different magnitudes of adversarial attacks.
-Performance with diverse network architecture.
-Dataset: CIFAR-10 ResNet-50 VGG-16 Inception-V3
-[…]
-MART (Wang et al. 2020) 75.05 49.05 4.57 66.72 44.27 3.60 76.05 49.59 5.91
-AWP (Wu, Xia, and Wang 2020) 75.59 50.80 3.91 67.14 45.97 3.47 76.74 50.46 4.07
-FSR (Kim et al. 2023) 78.38 50.59 3.97 69.59 44.90 3.30 77.28 50.02 4.73
-CFA (Wei et al. 2023) 77.88 50.64 4.27 70.45 45.53 4.66 76.70 49.97 4.46
-UIAT (Dong et al. 2023) 78.42 51.00 2.99 68.46 45.27 3.57 75.19 51.23 3.70
-SGLR (Li et al. 2024b) 80.39 50.55 5.21 70.90 45.04 3.94 80.62 50.73 5.00
-DHAT (Ours) 79.94 51.92 1.85 70.01 46.40 1.34 77.85 51.81 1.92
-DHAT-AWP (Ours) 76.84 52.05 2.34 67.41 47.21 2.18 77.12 52.08 2.88
-DHAT-CFA (Ours) 80.97 52.38 2.09 70.52 47.83 2.21 78.52 52.67 2.44
-Table 2: Comparison of robustness (%) and robust generalization gap (%). The number in bold indicates the best performance.
+4.3.1 Performance with diverse network architecture.
 To evaluate the adaptability of our method across different network architectures, we applied DHAT on ResNet-50, VGG-16, and Inception-V3.
@@
 Figure 3: Comparisons of robustness (%) under varying $\epsilon$ adversarial attacks using ResNet-18 on CIFAR-10.
-Different attack magnitude.
+4.3.2 Different attack magnitude.
 Figure 3 illustrates the performance of DHAT under varying magnitudes of PGD-10 and C&W attacks.
-Our findings reveal that DHAT maintains a significant performance advantage across all tested attack magnitudes.
+Our findings reveal that DHAT maintains a significant performance advantage across attack magnitudes.
 Compared to other methods, DHAT exhibits improved stability and robustness, with minimal performance degradation even under more severe attacks.
 This consistent performance across varying attack levels underscores the effectiveness of our approach in mitigating the impact of adversarial perturbations.
-Further details of the attack analysis are provided in the supplementary material.
-Ablation Analysis
+Further details are provided in the supplementary material (Section 7).
+Table 3: Comparison of computational costs (# A100 GPU hours) using different adversarial training methods on CIFAR-10.
+Method Architecture
+ResNet-18 WRN28-10 Inception-V3
+MART [45] 0.050 0.285 0.331
+AWP [49] 0.048 0.274 0.318
+FSR [21] 0.051 0.291 0.337
+CFA [47] 0.045 0.262 0.303
+UIAT [9] 0.040 0.245 0.283
+SGLR [27] 0.039 0.240 0.278
+DHAT (Ours) 0.041 0.248 0.287
+4.4 Computational Efficiency
+The computational costs of our method and baseline models are provided in terms of per-epoch training time on a single NVIDIA A100 GPU with a batch size of 128 on the CIFAR-10 dataset, as shown in Table 3.
+Our method demonstrates significantly improved time efficiency compared to the baseline MART, AWP, FSR, and CFA.
+While it requires slightly more computational resources than the state-of-the-art methods UIAT and SGLR, this minor increase is justifiable given the performance gains achieved.
+4.5 Ablation Analysis
 In this part, we conduct a comprehensive analysis of the impact of individual components and parameters of our model.
-Impact of individual components.
+4.5.1 Impact of individual components.
+Table 4: Comparison of robustness (%) and robust generalization gap (%). The number in bold indicates the best performance.
 […]
@@
 79.28 57.21 50.25 0.81
-Table 3: Comparison of robustness (%) and robust generalization gap (%). The number in bold indicates the best performance.
 We evaluate the contributions of the individual components of DHAT, specifically Debiased High-Confidence Logit Regularization (DHLR) and Foreground Logit Orthogonal Enhancement (FLOE), as shown in Table 4.
@@
 Figure 4: Analysis of hyper-parameters $\lambda_{1}$ and $\lambda_{2}$ using ResNet-18 on CIFAR-10. The robust accuracy denotes the accuracy under PGD-10 attack.
-Analysis of parameter sensitivity.
+4.5.2 Analysis of parameter sensitivity.
 Figure 4 illustrates the sensitivity of the optimization parameters $\lambda_{1}$ and $\lambda_{2}$ in DHAT.
@@
 Figure 5: Analysis of hyper-parameters $\omega$ using ResNet-18 on CIFAR-10 and Imagenette. The robust accuracy denotes the accuracy under PGD-10 attack.
-Analysis of threshold factor for quantifying spurious bias.
+4.5.3 Analysis of threshold factor for quantifying spurious bias.
 Figure 5 illustrates the impact of threshold values $w$ on datasets with varying resolutions.
@@
 For low-resolution datasets, the optimal threshold tends to be higher compared to high-resolution datasets, as high-resolution images more readily separate background features.
-Figure 6: Visualization of feature activation maps using UIAT and our proposed method DHAT, utilizing Grad-CAM under inverse attack ($\epsilon=-8$) and adversarial attack ($\epsilon=8$).
+Figure 6: Visualization of feature activation maps using UIAT and DHAT, utilizing Grad-CAM under inverse attack ($\epsilon=-8$) and adversarial attack ($\epsilon=8$).
 Left-to-right: the original image; normal feature activation maps; feature activation maps under adversarial attacks for UIAT and DHAT; and feature activation maps under inverse adversarial attacks for UIAT and DHAT.
-Discussion DHAT Effectiveness
+4.6 Discussion DHAT Effectiveness
 In this part, we analyze the mechanisms underlying the effectiveness of our proposed method, Debiased High-Confidence Adversarial Training (DHAT), with a focus on its capacity to mitigate spurious correlation bias.
-We begin by visualizing the feature activation maps of models trained with UIAT and DHAT under both adversarial attacks and inverse adversarial attacks.
-These visualizations, as depicted in Figure 6, highlight notable differences in feature attention maps between the models.
-Specifically, our proposed method DHAT restores the model’s attention to its normal state.
-In contrast, UIAT exhibits biased feature activation, particularly under inverse adversarial attack conditions, where the model’s attention shifts toward background features, reducing its attention to foreground features.
-DHAT’s effectiveness is attributed to its ability to recalibrate biased high-confidence logits during inverse adversarial attacks and reduce the correlation between high-confidence logits and background features in affine space.
-This ensures that the alignment with inverse adversarial examples is not compromised by spurious correlation bias.
+4.6.1 Feature Activation Map Analysis
+Figure 6 visualizes feature activation maps of UIAT and DHAT models under adversarial and inverse adversarial attacks.
+DHAT successfully maintains normal attention patterns, while UIAT exhibits biased activation with attention shifting significantly to background features during inverse attacks.
+DHAT achieves this robust performance by recalibrating biased high-confidence logits and systematically reducing correlation between logits and background features.
 Refer to caption Refer to caption
-(a) IoU with True Foreground (b) IoU with True Background
-Figure 7: Intersection over Union (IoU) of attention regions with true foreground and background regions. The comparison is conducted among standard adversarial training (AT), UIAT, and DHAT.
-Furthermore, we quantitatively evaluate DHAT’s impact on mitigating spurious correlation bias, we compute the Intersection over Union (IoU) between the model’s attention maps and the true foreground and background regions.
-Figure 7 illustrates the comparison under inverse adversarial attacks using conventional AT, UIAT, and DHAT models.
-Our results show that DHAT substantially improves the model’s attention to foreground features while effectively alleviating biased feature activation.
-Specifically, DHAT achieves higher IoU scores with true foreground regions and lower IoU scores with background regions compared to both AT and UIAT.
-These findings confirm that DHAT effectively mitigates spurious correlation bias.
-Conclusion
+(a) IoU with Normal Foreground (b) IoU with Normal Background
+Figure 7: IoU of attention regions with normal foreground and background regions. The comparison is conducted among conventional AT, UIAT, and DHAT on the Imagenette Dataset.
+Table 5: Comparison of training with different magnitude reverse adversarial samples (UIAT) and our debiased training DHAT using WRN28-10 on the CIFAR-10.
+[…]
+[…]
+UIAT -4 90.53 61.45 82.94 52.07 9.38
+UIAT -8 92.57 62.87 82.26 51.33 11.54
+UIAT -12 94.90 65.28 81.72 50.85 14.43
+DHAT (ours) -4 88.73 56.61 83.95 53.10 3.51
+4.6.2 Quantitative Evaluation of Bias Mitigation
+Figure 7 quantifies DHAT’s effectiveness through IoU between attention maps and true foreground/background regions.
+DHAT achieves higher foreground IoU and lower background IoU than both AT and UIAT, confirming its superior feature attention.
+Table 5 shows stronger inverse attacks in UIAT worsen overfitting, increasing the robust gap from 9.38 to 14.43.
+In contrast, DHAT achieves better test robustness (83.95% PGD-10, 53.10% AA) with a significantly smaller robust gap (3.51), effectively mitigating overfitting while maintaining strong performance.
+5 Conclusion
 In this work, we identify a critical issue in adversarial training: AT model exhibits biased feature activation under inverse adversarial attacks.
-Specifically, training with inverse adversarial examples leads to a shift in the model’s attention toward background features, resulting in spurious correlation bias.
+Specifically, training with inverse adversarial examples causes the model attention to shift from the foreground to background features, resulting in spurious correlation bias.
 To tackle this challenge, we propose a novel method called Debiased High-Confidence Adversarial Training (DHAT).
@@
 This work highlights DHAT’s effectiveness in enhancing model resilience and presents an adversarial training framework that integrates seamlessly with advanced training strategies.
-Supplementary
-This supplementary material provides an in-depth examination of the methodologies and results that support the experiments detailed in our study.
-It aims to offer comprehensive information that validates the findings and ensures reproducibility and transparency.
-Experimentation Details
+Acknowledgment
+This work is supported by the National Natural Science Foundation of China under Grant No.
+62276221, 62376232; Key Laboratory of Equipment Data Security and Guarantee Technology, Ministry of Education under Grant No.
+2024020200.
+Supplementary Materials
+This supplementary material offers a detailed examination of the methodologies and results that underpin the experiments in our study.
+It is designed to provide comprehensive information to validate the findings and ensure reproducibility and transparency.
+The content is organized as follows:
+  • Section 6: A comprehensive and detailed description of the experimental setup and methodology.
+  • Section 7: A formal presentation of the proposed algorithm, including its pseudocode representation.
+  • Section 8: In-depth evaluation of model performance under various transfer attack scenarios.
+  • Section 9: Detailed results from adversarial training using synthetic data for robustness assessment.
+  • Section 10: Extensive exploration of techniques for generating foreground-background attention maps.
+  • Section 11: Thorough examination of the model’s robustness under varying levels of attack intensity.
+6 Experimentation Details
 In this part, we provide a comprehensive overview of the experimental setup.
-Datasets Information
-CIFAR-10 (Krizhevsky, Hinton et al.
-2009) comprises 10 classes, with each image having a resolution of 32x32 pixels.
-This dataset is commonly used for basic image classification tasks and serves as a benchmark for evaluating the performance of various models on a relatively simple dataset.
-CIFAR-100 (Krizhevsky, Hinton et al.
-2009) contains 100 classes, with each image also having a resolution of 32x32 pixels.
-The expanded class set in CIFAR-100 increases complexity, enabling a more rigorous evaluation of models in distinguishing among a broader range of objects.
-Imagenette (Howard 2019), a subset of ImageNet, includes 10 classes with images at a higher resolution of 320x320 pixels.
-This dataset is constructed to provide a more realistic and challenging benchmark due to its increased resolution and closer representation of real-world images.
-Assessing models on datasets with a larger number of classes and higher resolutions, such as CIFAR-100 and Imagenette, facilitates in-depth analysis of their generalization capabilities.
-This ensures that models demonstrate effectiveness not only on simpler tasks but also exhibit resilience when confronted with more complex and diverse data.
-Training Setup
-In the training setup, we employ the Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9 and a weight decay factor of $5\times 10^{-4}$.
-The initial learning rate is set to 0.1 for CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al.
-2009) datasets, while 0.01 for Imagenette (Howard 2019).
-All models undergo 100 epochs of training, with the learning rate reduced by a factor of 0.1 at the 80th and 90th epochs.
-The trade-off hyper-parameters $\lambda_{1}$ and $\lambda_{2}$ are set to 1.0 and 1.0.
-To maintain experimental fairness, all comparative methods undergo identical training strategy, involving the same number of epochs and the same optimizer.
+6.1 Training Setup
+6.1.1 CIFAR-10 and CIFAR-100
+For the CIFAR-10 and CIFAR-100 [23] datasets, we employ the Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9 and a weight decay factor of $5\times 10^{-4}$.
+The initial learning rate is set to 0.1, and models undergo 100 epochs of training with the learning rate reduced by a factor of 0.1 at the 80th and 90th epochs.
+For adversarial training on these datasets, we create adversarial examples through a 10-iteration attack where the maximum $\ell_{\infty}$-norm of the adversarial perturbation is limited to $\epsilon=8/255$, using a step size of $\alpha=2/255$.
+To ensure reliability and fairness, we align the inverse adversarial example generation process with the UIAT settings [9], employing the same loss function and perturbation constrained by $\epsilon=4/255$.
+6.1.2 ImageNet-1K
+For the ImageNet-1K dataset, we strictly follow the training protocol established in [39] to ensure fair comparison with existing methods.
+Specifically, we implement a 2-iteration PGD over 50 epochs, and set $\epsilon=4/255$ and $\alpha=1/255$ for adversarial perturbations.
+This adherence to established benchmarking parameters enables direct and meaningful comparison with state-of-the-art approaches evaluated on this challenging large-scale dataset.
+6.1.3 Common Settings
+The hyperparameters $\lambda_{1}$ and $\lambda_{2}$ are consistently set to 1.0 across all datasets.
+To maintain experimental fairness, all comparative methods undergo identical training strategies with the same number of epochs and optimizer configurations.
 For Grad-CAM to quantify the spurious correlation bias, we use a pre-trained ResNet-18 model under simple training, introducing minimal additional training cost.
-The experiments were conducted on a single NVIDIA Tesla A100.
-For adversarial training, we create adversarial examples through a 10-iteration attack where the maximum $\ell_{\infty}$-norm of the adversarial perturbation is limited to $\epsilon=8/255$, using a step size of $\alpha=2/255$.
-To ensure the reliability of our experiments, we align the inverse adversarial example generation process with UIAT settings (Dong et al.
-2023), employing the same loss function and same perturbation constrained by $\epsilon=4/255$.
-Notably, our method does not employ label smoothing techniques.
-This choice is deliberate to avoid potential conflicts with efforts to mitigate spurious correlation bias, as label smoothing can blur class boundaries.
-Additionally, unlike UIAT, our approach does not utilize momentum terms to stabilize the generation process of inverse adversarial examples, focusing solely on the intrinsic properties of our method.
+All experiments were conducted on a single NVIDIA Tesla A100.
+Notably, our method does not employ label smoothing techniques to avoid potential conflicts with efforts to mitigate spurious correlation bias, as label smoothing can blur class boundaries and undermine the precision required for robust decision-making.
+Additionally, unlike UIAT, our approach does not utilize momentum terms to stabilize the generation process of inverse adversarial examples, instead focusing solely on the intrinsic properties and underlying dynamics of our method.
+Comprehensive training details are available in the supplementary material.
+6.2 Evaluation Setup
+Our evaluation consists of two main components: robustness performance and robust generalization.
+To evaluate robustness performance, we use PGD [29], C&W [4], and AutoAttack (AA) [7] within the $\ell_{\infty}$-norm.
+AutoAttack includes several attack methods such as APGD-DLR [7], APGD-CE [7], FAB [8], and Square [2].
+Adversarial attacks are generated using a step size of $\alpha=2/255$ and a specified maximum $\ell_{\infty}$-norm.
+Note that “Clean” indicates natural examples unaffected by adversarial perturbations.
+To evaluate robust generalization, we introduce the concept of the robust generalization gap [51] denoted as the “Robust Gap” quantifying the difference in robust performance between training and test sets under adversarial attacks.
+A smaller robust gap indicates improved robust generalization, reflecting reduced vulnerability to robust overfitting in the adversarial training model.
+7 Algorithm Details
+Algorithm 1 Debiased High-Confidence Adversarial Training (DHAT)
+0:  Training dataset $(X,Y)=\{(x_{i},y_{i})\}_{i=1}^{N}$; Network parameters $\theta$; Perturbation budget $\epsilon$; Hyperparameters $\lambda_{1},\lambda_{2}$; Attention threshold $\omega$.
+0:  Robust model parameters $\theta^{*}$
+1:  Initialize network parameters $\theta$
+2:  for each epoch do
+3:     for each mini-batch $(x,y)$ do
+4:        // Generate adversarial examples
+5:        $\hat{x}\leftarrow\arg\max_{||x^{\prime}-x||_{p}\leq\epsilon}\mathcal{L}_{AT}(f_{\theta}(x^{\prime}),y)$
+6:        $\hat{z}\leftarrow f_{\theta}(\hat{x})$
+7:        // Generate inverse adversarial examples
+8:        $\check{x}\leftarrow\arg\min_{||x^{\prime}-x||_{p}\leq\epsilon}\mathcal{L}_{Inv}(f_{\theta}(x^{\prime}),y)$
+9:        $\check{z}\leftarrow f_{\theta}(\check{x})$
+[…]
+11:        $M\leftarrow\mathcal{A}(x)$Different attention map computation techniques}
+12:        // Extract background features from inverse adversarial examples
+13:        $[\check{x}_{(B)}]_{(i,j)}\leftarrow\mathbb{I}_{(M_{i,j}<\omega)}\cdot\check{x}_{(i,j)}$
+14:        $\check{z}_{(B)}\leftarrow f_{\theta}(\check{x}_{(B)})$
+15:        // Compute debiased high-confidence logits
+16:        $\check{z}^{*}\leftarrow\check{z}-\check{z}_{(B)}$
+17:        // Compute DHLR loss
+18:        $\mathcal{L}_{DHLR}\leftarrow\mathcal{L}_{KL}(\phi(\check{z}^{*})||\phi(\hat{z}))$
+19:        // Compute FLOE loss
+20:        $\mathcal{L}_{FLOE}\leftarrow-|\check{z}-\frac{\check{z}\cdot\check{z}_{(B)}}{|\check{z}_{(B)}|^{2}}\cdot\check{z}_{(B)}|_{p}$
+21:        // Compute total loss
+22:        $\mathcal{L}_{DHAT}\leftarrow\mathcal{L}_{AT}(\hat{z},y)+\lambda_{1}\cdot\mathcal{L}_{DHLR}+\lambda_{2}\cdot\mathcal{L}_{FLOE}$
+23:        // Update network parameters
+24:        $\theta^{*}\leftarrow\theta-\nabla_{\theta}\mathcal{L}_{DHAT}$
+25:     end for
+26:  end for
+27:  return  $\theta^{*}$
+Algorithm 1 outlines the complete workflow of our proposed Debiased High-Confidence Adversarial Training (DHAT) framework, which systematically mitigates spurious correlations in adversarial training through two primary components: Debiased High-Confidence Logit Regularization (DHLR) and Foreground Logit Orthogonal Enhancement (FLOE).
+To efficiently compute attention maps, we employ Grad-CAM (Line 11) due to its computational efficiency.
+However, our framework is compatible with alternative saliency-based methods, such as SAM [22], which can provide enhanced performance at a higher computational cost.
+The extraction of background features (Lines 13-14) is performed using an adaptive threshold $\omega$, ensuring the identification of non-discriminative regions that contribute to spurious correlations.
+The overall training objective (Line 23) integrates standard adversarial training with our proposed debiasing regularization terms, weighted by $\lambda_{1}$ and $\lambda_{2}$.
+Despite incorporating additional minimal computational steps, the DHAT framework introduces only a marginal increase in resource consumption compared to standard adversarial training while significantly enhancing model robustness.
+Crucially, DHAT achieves these improvements without compromising clean-data accuracy, making it a highly practical and scalable solution for real-world adversarial defense applications.
+Table 6: Transfer attack accuracy (%) in the single-model transfer scenario. The number in bold indicates the best accuracy.
 Attack ($\epsilon=8$) Performance of UIAT / DHAT
@@
 AA 56.31/57.94 56.02/56.98 59.65/60.77 67.18/68.39 71.86/73.00 64.39/65.83 62.36/63.19 65.19/66.07 60.69/61.94
-Table 4: Transfer attack accuracy (%) in the single-model transfer scenario. The number in bold indicates the best accuracy.
-Refer to caption Refer to caption Refer to caption Refer to caption
-(a) FGSM Attack (b) PGD-10 Attack (c) PGD-20 Attack (d) C&W Attack
-Figure 8: Comparisons with varying $\epsilon$ values using ResNet-18 on the CIFAR-10. The $x$-axis represents the $\epsilon$ value, while $y$-axis represents the robust accuracy (%).
-Evaluation Setup
-Our evaluation consists of two main components: robustness performance and robust generalization assessment.
-To evaluate robustness performance, we use PGD (Madry et al.
-2018), C&W (Carlini and Wagner 2017), and AutoAttack (AA) (Croce and Hein 2020b) within the $\ell_{\infty}$-norm.
-AutoAttack includes several attack methods such as APGD-DLR (Croce and Hein 2020b), APGD-CE (Croce and Hein 2020b), FAB (Croce and Hein 2020a), and Square (Andriushchenko et al.
-2020).
-Adversarial attacks are generated using a step size of $\alpha=2/255$ and a specified maximum $\ell_{\infty}$-norm.
-Note that the “Clean” indicates natural examples unaffected by adversarial perturbations.
-To evaluate robust generalization, we introduce the concept of the robust generalization gap (Yin et al.
-2023) denoted as the “Robust Gap” quantifying the difference in robust performance between training and test sets under adversarial attacks.
-A smaller robust gap indicates improved robust generalization, reflecting reduced vulnerability to robust overfitting in the adversarial training model.
-Refer to caption Refer to caption Refer to caption Refer to caption
-(a) MART (b) UIAT (c) SGLR (d) DHAT
-Figure 9: Two-dimensional t-SNE embeddings of depth features extracted from the penultimate layer of the ResNet-18 model, trained on CIFAR-10 using various methods.
-The visualization illustrates the distinctiveness and compactness of the learned feature representations for each method.
-Performance Under Transfer Attack
-In this part, we evaluate our proposed model’s performance under transfer attacks and compare it with the UIAT (Dong et al.
-2023).
+8 Performance Under Transfer Attack
+In this part, we evaluate our proposed model’s performance under transfer attacks and compare it with the UIAT [9].
 Transfer attacks evaluate a model’s robustness by testing its performance against adversarial examples generated from different source models, which simulate real-world conditions where attackers may use varied strategies.
@@
 The table compares the performance of the UIAT method with that of our proposed model across different target models and transfer attack types.
-Robustness Performance
+8.1 Robustness Performance
 Our model consistently outperforms UIAT across all transfer adversarial attacks and target models.
@@
 This demonstrates the model’s robustness, highlighting its effectiveness not only in specific attack scenarios but also across various model architectures.
-Robustness Across Various Attack Types
+8.2 Robustness Across Various Attack Types
 Our proposed model DHAT demonstrates enhanced robustness, particularly against more challenging attacks such as PGD-50 and C&W.
@@
 This highlights the model’s superior capability to withstand adversarial perturbations.
-Generalization Across Various Source Models
+8.3 Generalization Across Various Source Models
 Our model exhibits strong performance across a range of source models, including VGG-16, WRN28-10, and ResNet-18.
 This indicates that the improved robustness of our model is not limited to specific attacks generated from source architectures but generalizes effectively across different adversarial settings.
-Robustness under Various $\epsilon$ Attacks
+Table 7: Comparison of robustness (%) and robust generalization gap (%) for models trained on generated data. The bolded numbers indicate the best performance.
+CIFAR-10 ResNet-18 WRN28-10
+[…]
+MART [45] 83.45 49.45 2.91 84.26 51.95 8.73
+AWP [49] 83.78 50.79 2.03 84.10 53.29 6.11
+FSR [21] 83.19 50.53 2.35 83.88 53.03 7.05
+CFA [47] 84.97 50.85 2.98 85.81 53.35 8.94
+UIAT [9] 85.10 52.09 3.04 86.73 54.59 9.41
+SGLR [27] 86.35 51.27 2.75 87.72 53.77 11.25
+DHAT (Ours) 87.62 54.42 0.83 88.94 56.92 2.48
+CIFAR-100 ResNet-18 WRN28-10
+[…]
+MART [45] 54.73 27.70 2.88 55.87 30.20 8.86
+AWP [49] 57.55 29.33 2.37 59.71 31.83 7.13
+FSR [21] 58.10 28.94 2.47 59.03 30.44 7.40
+CFA [47] 60.13 28.85 3.01 61.56 29.61 9.13
+UIAT [9] 59.92 28.48 4.47 60.24 30.98 14.32
+SGLR [27] 61.25 29.10 4.15 62.39 30.15 17.05
+DHAT (Ours) 63.14 32.21 0.98 64.80 34.71 2.96
+9 Performance with Generated Data
+We employ Diffusion Denoising Probabilistic Models (DDPM) [17] to generate an additional 50K samples for both CIFAR-10 and CIFAR-100 datasets following the [46].
+This synthetic data is then used to augment the training of both our method and all baseline models.
+The performance comparisons, as shown in Table 7, illustrate the impact of additional data on robustness and generalization.
+9.1 Robustness Performance
+The results in Table 7 demonstrate that our method outperforms baseline methods in terms of robustness when trained with the additional synthetic data.
+This improvement highlights the effectiveness of our approach in leveraging extra data to enhance model robustness, particularly under adversarial conditions.
+9.2 Generalization Performance
+We observe that, compared to the results in Table 1, most models trained with the additional data exhibit a reduced Robust Gap, indicating improved generalization.
+However, both UIAT and SGLR exhibit limited improvements in robust generalization.
+These methods rely on spurious correlations during training, which hinders their generalization performance even with the added data.
+Although UIAT and SGLR benefit from enhanced robustness due to increased data diversity, their robust generalization remains suboptimal, likely due to their dependence on non-essential features.
+This finding underscores the unique advantage of our method in achieving both robust accuracy and generalization with enriched datasets.
+Table 8: Comparison of robustness (%) and robust generalization gap (%) for using various attention map generation techniques using WRN28-10 on the CIFAR-10.
+[…]
+- 82.94 58.66 54.11 52.17 7.92
+Grad-CAM [33] 83.95 60.49 55.27 53.10 3.51
+Integrated-Grad [41] 83.97 60.35 55.18 53.04 3.68
+SOLO [44] 84.26 61.60 56.74 54.93 3.09
+SAM [22] 85.65 62.44 58.10 56.38 2.46
+10 Impact of Foreground-Background Recognition Techniques
+In this part, we investigate the influence of different attention map generation techniques on the performance of Debiased High-Confidence Logit Regularization (DHAT), as seen in Table 8.
+Our exploration extends beyond the initially employed Grad-CAM method to encompass a variety of attention map generation approaches, thereby providing a comprehensive analysis of their effects on model robustness and generalization.
+Refer to caption Refer to caption Refer to caption Refer to caption
+(a) FGSM Attack (b) PGD-10 Attack (c) PGD-20 Attack (d) C&W Attack
+Figure 8: Comparisons with varying $\epsilon$ values using ResNet-18 on the CIFAR-10. The $x$-axis represents the $\epsilon$ value, while $y$-axis represents the robust accuracy (%).
+10.1 Exploration of Various Attention Map
+The primary approach detailed in the main text utilizes Grad-CAM [33], a widely adopted method in weakly supervised object segmentation [28], to extract foreground and background feature maps efficiently.
+Grad-CAM offers a balance between computational efficiency and effectiveness, making it suitable for large-scale evaluations.
+However, to understand the broader applicability and potential improvements, we incorporated additional attention map generation techniques, including Integrated Gradients [41], SOLO [44], and SAM [22].
+10.2 Evaluation of Advanced Attention Map Models
+Incorporating more sophisticated attention map generation models, such as SAM, demonstrated enhanced robustness and reduced robust generalization gaps compared to simpler methods like Grad-CAM.
+These advanced models provide finer and more accurate segmentation of foreground and background features, which in turn leads to better alignment of high-confidence logits under adversarial conditions.
+However, this improvement in performance comes at the cost of increased computational overhead.
+More complex models require greater processing power and longer training times, which may be a limiting factor in resource-constrained environments.
+10.3 Trade-Off Between Performance and Computational Efficiency
+While the adoption of advanced attention map techniques can yield superior performance metrics, researchers must consider the trade-offs involved.
+Enhanced models may offer marginal gains in robustness and generalization, but the additional computational resources and time required may not always justify these benefits, especially in applications where real-time processing is essential.
+Therefore, the choice of attention map generation technique should be informed by the specific requirements and constraints of the deployment scenario.
+11 Robustness under Various Magnitude Attacks
 In this part, we evaluate the robustness of our proposed model under various levels of adversarial perturbations, quantified by different $\epsilon$ values.
-We compare our model with two baseline methods, MART (Wang et al.
-2020) and UIAT (Dong et al.
-2023), and the state-of-the-art SGLR method (Li et al.
-2024b) across various attack types, including FGSM, PGD-10, PGD-20, and C&W.
+We compare our model with two baseline methods, MART [45] and UIAT [9], and the state-of-the-art SGLR method [27] across various attack types, including FGSM, PGD-10, PGD-20, and C&W.
 We aim to demonstrate the superior generalization and robustness of our model under progressively challenging adversarial conditions.
@@
 The comparative analysis of our model with MART, UIAT, and SGLR reveals several key insights:
-Robustness Across Various Attack Types
+11.1 Robustness Across Various Attack Types
 Our model consistently outperforms the baselines and SGLR across all attack types and $\epsilon$ values.
 This consistent superiority underscores our model’s robust generalization ability, allowing it to maintain high accuracy even under more challenging adversarial attack conditions.
-Generalization Against Various Attacks
+11.2 Generalization Against Various Attacks
 The results suggest that the defensive mechanisms in our model are highly effective in mitigating the impact of various adversarial attacks (i.e., FGSM, PGD-10, PGD-20, C&W).
 This robustness is especially evident in scenarios with stronger attacks and higher $\epsilon$ values, where our model shows superior performance compared to existing methods.
-Practical Applicability
+11.3 Practical Applicability
 The enhanced robustness of our model across different $\epsilon$ values and attack types demonstrates its practical applicability in real-world scenarios, where adversarial perturbations can vary in strength and sophistication.
-Visualization of Learned Feature Distributions
-In this part, we provide a comparative analysis of the feature distributions learned by our proposed method, DHAT, and benchmark it against MART (Wang et al.
-2020), UIAT (Dong et al.
-2023), and SGLR (Li et al.
-2024b).
-Figure 9 illustrates these feature distributions using t-SNE visualization, with embeddings obtained from the penultimate layer of the ResNet-18 model trained on CIFAR-10.
-Our method, DHAT, demonstrates a marked improvement in feature representation.
-Specifically, the features learned by DHAT are more compact and distinct compared to those produced by the other methods.
-This enhanced discriminability stems from DHAT’s effectiveness in mitigating spurious correlation bias.
-By aligning adversarial examples with debiased high-confidence logits, DHAT facilitates the learning of features that are both robust and discriminative, leading to improved performance, particularly in margin-based attacks such as C&W attacks.
-When compared between MART and UIAT, MART demonstrates superior feature distinguishability.
-This advantage stems from MART’s alignment of logit distributions based on natural examples, which promotes the learning of more distinguishable features.
-In contrast, UIAT aligns the logits of adversarial examples with those of inverse adversarial examples, which are correlated with biased feature activation, resulting in less distinguishable feature representations.
-Additionally, SGLR’s performance in feature distinguishability is inferior to that of DHAT.
-This is because SGLR relies on soft-label distributions during training, which tends to rely on spurious correlations, thereby impairing class differentiation.
-In summary, DHAT demonstrates superior performance in t-SNE feature distributions, showcasing enhanced feature distinguishability and more effective representations.
-This improvement is attributed to DHAT’s ability to mitigate spurious correlation bias, which facilitates more robust and accurate feature representations.
-Refer to caption Refer to caption
-(a) UIAT (b) DHAT
-Figure 10: The learning curves for natural accuracy (%) and adversarial accuracy (%) under PGD-10 attack of ResNet-18 on CIFAR-10 dataset.
-Robust Overfitting Analysis
-In this part, we compare the learning curves of DHAT and UIAT to analyze robust overfitting and generalization.
-Experiments were conducted using ResNet-18 on the CIFAR-10 dataset, with a total of 200 training epochs for better visualization.
-Learning rate decay was applied at epochs 100 and 150, reducing the learning rate to 0.1.
-Figure 10 shows the learning curves for DHAT and UIAT.
-Our analysis shows that neither DHAT nor UIAT exhibits a significant or irreversible decline in robustness performance after learning rate adjustments (Yu et al.
-2022).
-This indicates that both methods maintain consistent robustness throughout the training process.
-However, DHAT consistently shows a smaller robust generalization gap and standard generalization gap between the training and test sets than UIAT.
-This reduced performance gap underscores DHAT’s superior generalization ability across different distributions, such as adversarial attack conditions or transitions to unseen test sets.
-The reduced performance disparity implies that DHAT effectively combats robust overfitting, leading to more reliable and consistent adversarial robustness.
-In contrast, UIAT shows a more pronounced performance decrease, indicating the degraded generalization capability.
-In conclusion, these findings confirm that DHAT not only enhances adversarial robustness but also improves robust generalization, representing a significant advancement in adversarial training techniques.